### v1

# Normal Quaternionic Matrices and Finitely Generated Witt Rings

[…]
Normal Quaternionic Matrices and Finitely Generated Witt Rings
Nico Lorenz N. Lorenz, Fakultät für Mathematik, Ruhr-Universität Bochum, Universitätsstraße 150, 44780 Bochum, Germany nico.lorenz@ruhr-uni-bochum.de  and  Alexander Schönert A. Schönert, Fakultät für Mathematik, Technische Universität Dortmund, Vogelpothsweg 87, 44227 Dortmund Germany alexander.schoenert@tu-dortmund.de
(Date: May 20, 2025)
Key words and phrases:
Quadratic Forms, Abstract Witt Rings, Quaternion Algebras, Elementary Type Conjecture
1991 Mathematics Subject Classification:
11E81, 16K50
0. Introduction
In attempts to characterize (finitely) generated Witt rings over fields of characteristic different from two, there have been abstract approaches for more than 50 years now, starting with the work of Knebusch, Rosenberg, and Ware in the early 1970s ([KnebuschRosenbergWare_WittRings]). In our context, abstract means that we try to grasp the ring structure of Witt rings by mimicking the structure of the square class group along with the represented values of binary quadratic forms, leading to quadratic form schemes as introduced by Cordes in [Cordes_FiniteNoQuaternions], or the behavior of 2-fold Pfister forms and their associated quaternion algebras, leading to quaternionic structures as in Marshall’s book in 1980 [Marshall_AbstractWittRings]. The idea of looking at Witt rings from this perspective has its origin in the following result of Cordes from 1973. Note that we assume all fields to be of characteristic different from two whenever we do not explicitly consider the field $\mathbb{F}_{2}$ with cardinality $2$.
Proposition 0.1.
(Cordes [Cordes_WittGroupEquivalenve, 2.2/2.3])
Let $E,F$ be fields with $G_{E}:=E^{\ast}/E^{\ast 2}$ and $G_{F}:=F^{\ast}/F^{\ast 2}$. The Witt rings of $E$ and $F$ are isomorphic as rings if and only if there is a group isomorphism $\Phi:G_{E}\to G_{F}$ with $\Phi(-E^{\ast 2})=-F^{\ast 2}$ preserving the isomorphy type of quaternion algebras, that is,
$\forall a,b,c,d\in G_{E}:\quad(a,b)_{E}\cong(c,d)_{E}\iff(\Phi(a),\Phi(b))_{F}\cong(\Phi(c),\Phi(d))_{F}.$
We focus on the approach to describe Witt rings using the properties of quaternion algebras (resp. their norm forms which are $2$-fold Pfister forms) and mainly follow Marshall’s book [Marshall_AbstractWittRings] for our basic notations and definitions.
All so far known examples of finitely generated Witt rings are of the so called elementary type, meaning that they can be derived from certain Witt rings only using two simple construction types. This gives rise to the elementary type conjecture (cf. Remark 2.5) stating that in fact all finitely generated Witt rings are of such type. It is known that the elementary type conjecture holds for Witt rings and thus quaternionic structures over fields $F$ of small order $n:=\operatorname{log}_{2}(|F^{\ast}/F^{\ast 2}|)\in\mathbb{N}_{0}$. The case $n\leq 2$ can be taken care of trivially (cf. Example 1.2), while the case $n=3$ was first done by Kula, Szczepanik and Szymiczek [KulaSzczepanikSzymiczek_8SquareClasses] and the case $n\in\{4,5\}$ by Carson and Marshall [CarsonMarshall_Decomposition]. The case $n=4$ was also dealt with by Szczepanik in 1985 independently ([Szczepanik_RadicalNotExceeding16]). In this paper we introduce a new perspective involving an abstract version of the $2$-torsion part of the Brauer group and developing normal quaternionic matrices which uniquely display quaternionic structures. We apply this to prove some hitherto open cases for the elementary type conjecture, in particular showing that the elementary type conjecture holds for order up to $6$. We also verify the elementary type conjecture for order up to $7$ in all but a few cases in which our computations get too time consuming. While the theoretical parts from Section 1 to Section 5 were developed by the second author in terms of his PhD thesis, the computational approach described in Section 6 was obtained by a joint effort of the authors.
We now give a brief outline of the structure of the article at hand. In Section 1, we introduce our notation and give the definition of quaternionic structures. Section 2 will be used to collect all necessary ingredients to state the elementary type conjecture along with some counting arguments which will be needed later on. This will be followed up by Section 3 where an abstract analogue to the $2$-torsion part of the Brauer group is explored. In Section 4, we define so called quaternionic matrices as a means to efficiently compute and describe quaternionic structures of small order. We also introduce the already mentioned normal quaternionic matrices which are then further investigated in Section 5. Finally, in the last Section 6, we present the algorithm and the computational means by which we prove our main result.
1. Quaternionic Structures
We summarize the main definitions and results needed for this article from Marshall’s book [Marshall_AbstractWittRings]. The results of classical quadratic form theory mentioned here can, for example, be found in Lam’s standard textbook [Lam_IntroQF].
Throughout this paper, we often consider groups $G$ of exponent $2$, which means that every element in $G$ is its own inverse. These can be viewed as vector spaces over $\mathbb{F}_{2}$, so by the dimension of $G$ we always mean its dimension as an $\mathbb{F}_{2}$ vector space. This dimension is finite if and only if $G$ is finite and in this case, we have $|G|=2^{n}$ for some $n\in\mathbb{N}_{0}$.
Quaternionic structures as defined below imitate the properties of isomorphy types of quaternion algebras over a given field of characteristic different from $2$.
Definition.
Let $G$ be a multiplicative group of exponent $2$ with neutral element $1\in G$ and a distinguished element $-1\in G$ (the case $-1=1$ being possible). Write $-a:=(-1)a$ for $a\in G$. Let $Q$ be a set with distinguished element $0\in Q$ and $q:G\times G\to Q$ a surjective map. We call the elements of $Q$ quaternions. We say that $S:=(G,-1,q)$ is a quaternionic structure if the following hold for all $a,b,c,d\in G$:
  1. (Q1)
    $q(a,-a)=0$
  2. (Q2)
    $q(a,b)=q(b,a)$
  3. (Q3)
    $q(a,b)=q(a,c)\iff q(a,bc)=0$
  4. (Q4)
    $q(a,b)=q(c,d)\iff\exists x\in G:\ q(a,b)=q(a,x)=q(c,x)=q(c,d)$
Furthermore, we define the radical of $S$ as
$\operatorname{rad}(S):=\{a\in G\mid q(a,x)=0\ \forall x\in G\}.$
We say that $S$ is degenerate if $\operatorname{rad}(S)\not=\{1\}$, and else nondegenerate. If $G=\operatorname{rad}(S)$, we say that $S$ is totally degenerate.
If $S^{\prime}=(G^{\prime},-1^{\prime},q^{\prime})$ is another quaternionic structure, we say that $S$ and $S^{\prime}$ are isomorphic, in symbols $S\cong S^{\prime}$, if there exists a group isomorphism $\phi:G\to G^{\prime}$ satisfying $\phi(-1)=-1^{\prime}$ and for all $a,b\in G$:
$q(a,b)=0\iff q^{\prime}(\phi(a),\phi(b))=0.$
We refer to $\dim(G)$ as the order of $S$. Note that we sometimes refer to $G$ as the square class group (of $S$).
Remark 1.1.
  1. (a)
    Every field $F$ defines a quaternionic structure, using $G_{F}:=F^{\ast}/F^{\ast 2}$ and $q_{F}(a,b):=(a,b)_{F}$, the quaternion algebra with slots $a$ and $b$ over $F$ (for (Q4), see [Lam_IntroQF, Common Slot Theorem III.4.13]). Here, the zero-element in $Q$ can be viewed as the split quaternion algebra or the zero in the Witt ring. We write $S_{F}:=(G_{F},-F^{\ast 2},q_{F})$.
  2. (b)
    It is an easy exercise to show that quaternionic structures are isomorphic if and only if the abstract analogue of the condition from Proposition 0.1 is satisfied, justifying the above definition.
  3. (c)
    It is an immediate consequence of the axioms (Q1), (Q2) and (Q3) that any quaternionic structure $(G,-1,q)$ satisfies $q(a,1)=0$ and for all $a,b\in G$, we have:
    $q(a,b)=q(a,-ab)=q(b,-ab)$
We include some basic examples along with their notation:
Example 1.2.
  • There is only one quaternionic structure of order zero, denoted as $\mathbb{L}_{0}$, and defined by the unique map $\{1\}\times\{1\}\to\{0\}$. It is realized as $S_{F}$ for any quadratically closed field $F$ of characteristic different from $2$, for example $\mathbb{L}_{0}\cong S_{\mathbb{C}}$. This structure is both totally degenerate and nondegenerate by definition.
  • There are exactly two non isomorphic totally degenerate quaternionic structures of order $n\in\mathbb{N}$, depending on the choice whether or not $-1=1$ in the square class group $G$ and defined by the unique map $G\times G\to\{0\}$. For $n=1$, we denote this structure by $\mathbb{L}_{1,0}$ if $-1=1$ and $\mathbb{L}_{1,1}$ if $1$-1$\not$=1- 1 ≠ 1.
  • There are three quaternionic structures $S=(\{1,a\},-1,q)$ of order $1$: Firstly, if $-1=1$, then $q(1,1)=q(1,a)=q(a,1)=q(a,a)=0$, so $S\cong\mathbb{L}_{1,0}$. Secondly, if $-1=a$ and $S\not\cong\mathbb{L}_{1,1}$, then it can be easily verified that $S\cong S_{\mathbb{R}}=:\mathbb{L}_{1}$. Note that also $\mathbb{L}_{1,0}$ and $\mathbb{L}_{1,1}$ are realized by finite fields of cardinality $\equiv_{4}1$ resp. $\equiv_{4}3$.
  • Let $S=(G,-1,q)$ be a finite nondegenerate quaternionic structure of order $n\geq 3$ with set of quaternions $Q$ with $|Q|=2$, called a local (quaternionic) structure. Then $q$ can be viewed as a nondegenerate symmetric bilinear map $G\times G\to\mathbb{F}_{2}$ (recall that $G$ is an $\mathbb{F}_{2}$ vector space). Conversely, any such nondegenerate symmetric bilinear map defines a local structure. There are exactly two local structures for even $n$, written $\mathbb{L}_{n,0}$ (the case $-1=1$) and $\mathbb{L}_{n,1}$ (the case $1$-1$\not$=1- 1 ≠ 1). Also, there is only one local structure for any odd $n$, written $\mathbb{L}_{n}$. All the local quaternionic structures can be realised by certain field extensions of $\mathbb{Q}_{2}$, the $2$-adic completion of $\mathbb{Q}$, cf. [Marshall_AbstractWittRings, Chapter 5].
Remark 1.3.
We do not give the definition of abstract Witt rings here as we do not need it explicitly. We refer to [Marshall_AbstractWittRings, 4.5] where it is shown that the categories of quaternionic structures and abstract Witt rings as defined there are equivalent. We only outline the notation for quadratic forms and important properties that will be employed here. From a classical standpoint, the quadratic forms considered here correspond to nondegenerate and diagonalized quadratic forms in the classical sense:
  • A (quadratic) form over $S$ of dimension $n\in\mathbb{N}$ is an $n$-tuple, denoted $\langle a_{1},\ldots,a_{n}\rangle\in G^{n}$.
  • Isometry of quadratic forms (denoted by $\cong$) is defined inductively as in ([Lam_IntroQF, §I, Exercise 19]) and isotropy derived from it canonically. Isometry defines an equivalence relation (technical proof by Marshall in [Marshall_AbstractWittRings, 2.6]) and isometry is invariant under permutation of tuple entries which means that the following is well defined.
  • For quadratic forms $\varphi=\langle a_{1},\ldots,a_{n}\rangle$ and $\psi:=\langle b_{1},\ldots,b_{m}\rangle$, the sum and product are
    $\varphi\perp\psi:=\langle a_{1},\ldots,a_{n},b_{1},\ldots,b_{m}\rangle\quad\text{and}\quad\varphi\otimes\psi:=\perp_{i=1}^{m}b_{i}\varphi:=\perp_{i=1}^{m}\langle b_{i}a_{1},\ldots,b_{i}a_{n}\rangle.$
  • The hyperbolic plane $\mathbb{H}:=\langle 1,-1\rangle$ is the unique isotropic form of dimension $2$. A quadratic form is called hyperbolic if it is isometric to the $k$-fold sum of hyperbolic planes $k\times\mathbb{H}$ for some $k\in\mathbb{N}_{0}$.
  • Given $n\in\mathbb{N}_{0}$ and $a_{1},\ldots,a_{n}\in G$, we say that $\langle\!\langle a_{1},\ldots,a_{n}\rangle\!\rangle:=\bigotimes_{i=1}^{n}\langle 1,-a_{i}\rangle$ is an $n$-fold Pfister form.
  • The set of represented elements of a form $\varphi$ over $S$ is $D_{S}(\varphi):=\{a\in G\mid\varphi\cong\langle a,\ldots\rangle\}$.
  • The value group of $a\in G$ is defined as the set $V(a):=D_{S}(\langle\!\langle a\rangle\!\rangle)$ and can also be written as $\{b\in G\mid q(a,b)=0\}$. Note that in literature, similar to the ambiguity of the sign convention for Pfister forms, this set is sometimes defined as $V(-a)$.
  • Defining the (abstract) Witt ring $W(S)$ of a quaternionic structure $S=(G,-1,q)$ now works in most parts analogue to the classical case, by establishing Witt cancellation and Witt equivalence in the abstract setting.
  • Viewed as elements of the Witt ring of $S$, the $1$-fold Pfister forms additively generate the fundamental ideal $I(S)$ of (Witt equivalence classes of) even dimensional forms. As in the classical theory, the $k$-th power $I^{k}(S)$ of $I(S)$ is additively generated by the $k$-fold Pfister forms. Note that our notation usually does not distinguish between quadratic forms and their Witt equivalence class as an element of the Witt ring whenever the point of view is clear from the context.
Lastly, consider the following property for the $k$-th power of the fundamental ideal, with $k\in\mathbb{N}$:
($AP_{k}$) If φIk(S){0} is anisotropic, then $\text{If $\varphi\in I^{k}(S)\setminus\{0\}$ is anisotropic, then}\dim(\varphi)\geq 2^{k}$
In the case that $S$ comes from a field, the above holds for arbitrary $k\in\mathbb{N}$, which is known as the Arason-Pfister Hauptsatz [ArasonPfister_Krull]. Generally, $AP_{1}$ and $AP_{2}$ hold in abstract Witt rings (cf. [Marshall_AbstractWittRings, 3.23]) while it is an open question whether or not $AP_{k}$ holds for any $k\geq 3$. Especially $AP_{3}$ plays an important role considering the abstract Brauer group as can be seen later. For a quaternionic structure satisfying $AP_{3}$ we henceforth simply say $AP_{3}$ structure.
2. The Elementary Type Conjecture
In this section, we recollect all ingredients for the elementary type conjecture as stated by Marshall [Marshall_AbstractWittRings, Chapter 5, Problem 4].
Proposition 2.1.
Let $S=(G,-1,q)$, $S_{1}=(G_{1},e_{1},q_{1})$ and $S_{2}=(G_{2},e_{2},q_{2})$ be quaternionic structures with sets of quaternions $Q,Q_{1},Q_{2}$.
The direct product $S_{1}\times S_{2}:=(G_{1}\times G_{2},(e_{1},e_{2}),q_{1}\times q_{2})$ with quaternions $Q_{1}\times Q_{2}$
$q_{1}\times q_{2}((a_{1},a_{2}),(b_{1},b_{2})):=(q_{1}(a_{1},b_{1}),q_{2}(a_{2},b_{2}))$
and zero-quaternion $(0,0)$ is a quaternionic structure with abstract Witt ring
$W(S_{1}\times S_{2})=(W(S_{1})\times W(S_{2}))_{\text{par}}:=\{(\varphi,\psi)\mid\dim(\varphi)\equiv_{2}\dim(\psi)\}.$
Let $G_{x}$ be the direct product of $G$ with a cyclic group $\{1,x\}$ of order $2$, written as an inner direct product. The group extension $S[x]:=(G_{x},-1,q_{x})$ with quaternions in $Q\times G$,
$q_{x}(ax^{\alpha},bx^{\beta})=(q(a,b),(-1)^{\alpha\beta}a^{\beta}b^{\alpha})$
and zero quaternion $(0,1)$ is a quaternionic structure with Witt ring $W(S[x])=W(S)[\{1,x\}]$, the group ring of the Witt ring of $S$ and the cyclic group of order 2. We often write $S\Delta:=S[x]$. Constructing the group extension can be iterated and we write $S\Delta_{n}$ for the $n$-fold group extension of $S$ for any $n\in\mathbb{N}$.
Remark 2.2.
If in the above proposition, $S\cong S_{F}$ for a field $F$, then $S_{F}\Delta\cong S_{E}$ for the field $E:=F(\!(x)\!)$ of formal Laurent series over $F$. In 1985, applying quite some valuation theory, Kula [Kula_FieldsQFSchemes] showed that if $S_{1}$ and $S_{2}$ are realized by fields, then so is $S_{1}\times S_{2}$.
Example 2.3.
  • The totally degenerate quaternionic structure of order $n\in\mathbb{N}$ (cf. Example 1.2) are exactly $\mathbb{L}_{1,0}^{n}$ (the $n$-fold direct product) and $\mathbb{L}_{1,1}^{n}$.
  • We have $\mathbb{L}_{0}\Delta\cong\mathbb{L}_{1,0}$ which also follows from Example 1.2 as there is only one quaternionic structure of order $2$ and level $1$.
  • We have $\mathbb{L}_{1}\times\mathbb{L}_{1}\cong\mathbb{L}_{1}\Delta$. This is in fact the only decomposable (meaning that it can be written as the direct product of two structures of order at least 1) group extension (cf. [Marshall_ETC, p. 8]).
Marshall introduced the following normalized decomposition ([Marshall_AbstractWittRings, 5.12]).
Proposition 2.4.
Let $S$ be a finite quaternionic structure with $|\operatorname{rad}(S)|=2^{m}$. Choose $\epsilon=0$ if $-1=1$ in $S$ and $\epsilon=1$ else. There exist indecomposable nondegenerate quaternionic structures $S_{1},\ldots,S_{k}$ such that
$S\cong\mathbb{L}_{1,\epsilon}^{m}\times S_{1}\times\ldots\times S_{k}$
and the $S_{i}$ are unique up to isomorphy and permutation.
Definition.
Let $S$ be a finite quaternionic structure. We call $S$ of simple elementary type if $S$ is isomorphic to one of $\mathbb{L}_{0},\mathbb{L}_{1,0},\mathbb{L}_{1,1},\mathbb{L}_{1}$ or a local structure. We call $S$ of elementary type if $S$ can be obtained from simple elementary types applying only direct products and group extensions.
Remark 2.5.
The elementary type conjecture (ETC) states that all finite quaternionic structures are in fact of elementary type. Along with Kula’s result mentioned in Remark 2.2, it follows that all elementary type structures are realized by fields. Since it is not yet known whether all quaternionic structures are in fact realized by fields, a possibly weaker version of this conjecture is that all finite quaternionic structures which are realised by fields are of elementary type, this version sometimes being referred to as the weak elementary type conjecture (wETC). Both versions have so far only been proven for special cases. Some of them will be stated at the end of this section. However, we first give an overview of elementary type structures of small order, also considering some important invariants.
The following invariants of isomorphy types of quaternionic structures can be defined as in the classical setting:
Definition.
Let $S=(G,-1,q)$ be a quaternionic structure. We define the level of $S$ as
$s(S):=\operatorname{inf}\{k\in\mathbb{N}\mid-1\in D_{S}(k)\}$
and the Pythagoras number of $S$ as
$p(S):=\operatorname{inf}\{k\in\mathbb{N}\mid D_{S}(k)=D_{S}(k+1)\}$
where $\operatorname{inf}(\emptyset)=\infty$ and we say that $S$ is pythagorean if $p(S)=1$. Furthermore, we define the u-invariant of $S$ as
$u(S):=\operatorname{sup}\{\dim(\varphi)\mid\varphi\text{ is an anisotropic torsion form over }S\}$
where ’torsion form’ means that there exists $n\in\mathbb{N}$ such that $n\times\varphi$ is hyperbolic. In the classical sense, this is often referred to as the general u-invariant (cf. [Lam_IntroQF, Chapter XI]).
If $F$ is a field, note that the values $s(S_{F})$ and $p(S_{F})$ are the level and Pythagoras number of $F$. In particular, we point out that the level of $S$ is generally not the ring theoretic level of the associated Witt ring.
Example 2.6.
With Marshall’s normalized decomposition, it is possible to iteratively compute complete lists of the elementary types of a certain order. We give those for order up to 2, including the level $s$, Pythagoras number $p$, and u-invariant $u$ in Table 1. As they will be important later on, we also include the dimension $d_{1}:=\dim(V(-1))$ and the minimum $d_{\operatorname{min}}$ over all $\dim(V(a))$ for $a\in G$. We also include the general values for direct products, group extensions, and local structures so that all the given invariants can be computed for all elementary types with this list. Similar to Example 1.2, it can be computed in a straightforward manner that the list contains all quaternionic structures of order up to $2$. Take $k\in\mathbb{N}$, $k\geq 2$.
structure order $s$ $p$ $u$ $d_{1}$ $d_{\operatorname{min}}$
$\mathbb{L}_{0}$ 0 1 1 1 0 0
$\mathbb{L}_{1,0}$ 1 1 2 2 1 1
$\mathbb{L}_{1,1}$ 1 2 2 2 1 1
$\mathbb{L}_{1}$ 1 $\infty$ 1 0 0 0
$\mathbb{L}_{1,0}^{2}$ 2 1 2 2 2 2
$\mathbb{L}_{1,1}^{2}$ 2 2 2 2 2 2
$\mathbb{L}_{1,1}\times\mathbb{L}_{1}$ 2 $\infty$ 2 2 1 2
$\mathbb{L}_{1}^{2}\cong\mathbb{L}_{1}\Delta$ 2 $\infty$ 1 0 0 0
$\mathbb{L}_{1,0}\Delta$ 2 1 2 4 2 1
$\mathbb{L}_{1,1}\Delta$ 2 2 3 4 1 1
$\mathbb{L}_{2k-1}$ $12k-$ 1 4 4 4 $22k-$ 2 $22k-$ 2
$\mathbb{L}_{2k,0}$ $2k$ 1 2 4 $2k$ $12k-$ 1
$\mathbb{L}_{2k,1}$ $2k$ 2 3 4 $12k-$ 1 $12k-$ 1
$S$ $n$ $s$ $p$ $u$ $d_{1}$ $d_{\operatorname{min}}$
$S^{\prime}$ $n^{\prime}$ $s^{\prime}$ $p^{\prime}$ $u^{\prime}$ $d_{1}^{\prime}$ $d_{\operatorname{min}}^{\prime}$
$S\Delta$ $n+1$ $s$$\begin{cases}p\text{ if }s=\infty\\ s+1\text{ else}\end{cases}$ $2u$ $d_{1}$ 1
$S\times S^{\prime}$ $n+n^{\prime}$ $\operatorname{max}\{s,s^{\prime}\}$ $\operatorname{max}\{p,p^{\prime}\}$ $\operatorname{max}\{u,u^{\prime}\}$ $d_{1}+d_{1}^{\prime}$ $d_{\operatorname{min}}\!+\!d_{\operatorname{min}}^{\prime}$
Table 1. List of small and local structures
Remark 2.7.
There have been many attempts to find a field with finitely many square classes and finite level $\geq 8$. This gave rise to the level conjecture (LC) which states that all finite quaternionic structures (coming from fields) have level in $\{1,2,4,\infty\}$. It holds for elementary type structures and thus the LC would follow from the ETC. In fact, the previous example shows that every elementary type structure has level in $\{1,2,4,\infty\}$, Pythagoras number in $\{1,2,3,4,5\}$ and $u$-invariant in $\{0\}\cup\{2^{m}\mid m\in\mathbb{N}_{0}\}$ and all these values can be obtained (though not independently, since for example the Pythagoras number cannot exceed the level by more than $1$). The ’smallest’ example for a quaternionic structure with pythagoras number $5$ is $\mathbb{L}_{3}\Delta$. Note that the LC is known to hold for Witt rings (of fields) of order up to 8 by a result of Becher in 2001 [Becher_NoSquareClasses].
Definition.
Let $S=(G,-1,q)$ be a quaternionic structure and $k\in\mathbb{N}_{0}$. We say that $a\in G$ is $k$-rigid if $\dim(D_{S}(\langle 1,a\rangle))=k$. If $G$ contains a $k$-rigid element, we say that $S$ is $k$-rigid. If $k\geq 1$ and $G$ additionally does not contain a $(k-1)$-rigid element, we say that $S$ is strictly $k$-rigid. We usually only say rigid and strictly rigid instead of $1$-rigid and strictly $1$-rigid.
Note that $S$ is 00-rigid if and only if $1\in G$ is 00-rigid if and only if $S$ is pythagorean.
We state a few results on the ETC, where especially those regarding $k$-rigid elements for small $k$ will be important for reducing the length of the computations in Section 6.
Proposition 2.8.
Let $S$ be a finite quaternionic structure of order $n\in\mathbb{N}$.
  1. (a)
    Marshall in [Marshall_AbstractWittRings, 6.23]: If $S$ is pythagorean, then every indecomposable factor of $S$ is isomorphic to $\mathbb{L}_{1}$ or a group extension of $\mathbb{L}_{1}$. In particular, $S$ is of elementary type.
  2. (b)
    Kula in [Kula_FinitelyGeneratedWittRings, 3.12]: Every quaternionic structure with at most 12 quaternions is of elementary type.
  3. (c)
    Carson and Marshall in [CarsonMarshall_Decomposition, § 5]: If $n\leq 5$, then $S$ is of elementary type.
  4. (d)
    Assuming that every quaternionic structure of order up to $n-1$ is of elementary type, $S$ is also of elementary type in the following cases:
    • $S$ is degenerate (Marshall, cf. normalized decomposition in 2.4)
    • $S$ is nondegenerate and rigid (Bos in [Bos_StructureWittRing, 2.10])
    • $s(S)=1$ and $S$ is $2$-rigid (Carson and Marshall in [CarsonMarshall_Decomposition, 3.10])
    • $s(S)=2$ and there is a $2$-rigid element in $D_{S}(\langle 1,1\rangle)$ (Cordes in [Cordes_BinaryRepr4, 3.1/4.1])
    • $s(S)=4$ and $1$ is a $2$-rigid element (Cordes in [Cordes_BinaryRepr4, 2.6]).
To close this section, we include a technical result to count elementary types. This is a slightly refined version of the one given in ([CarsonMarshall_Decomposition, 4.5]).
Notation.
We write $\mathcal{E}_{k}(n)$ for the set of isomorphy tyes of strictly $k$-rigid nondegenerate elementary type structures of order $n$ and
¯$\overline{\mathcal{E}}_{k}(n):=\{S\in\mathcal{E}_{k}(n)\mid S\text{ is indecomposable}\}.$
[…]
Definition.
Let $k\in\mathbb{N}_{0}$ and $n\in\mathbb{N}$. We define the set $P_{k,n}$ as
[…]
Note that $P_{0,n}$ is the set of partitions of $n$.
Lemma 2.9.
Let $k,n\in\mathbb{N}_{0}$. The following hold:
  1. (a)
    $e_{\text{d}}(n)=e(n-1)+e^{\prime}(n-1)-e^{\prime}(n-2)$ for $n\geq 2$
  2. (b)
    $e_{\Delta}(n)=e(n-1)$ for $n>2$.
  3. (c)
    We have $e_{0}(0)=e_{0}(1)=e_{0}(2)=1=\overline{e}_{0}(1)$, $\overline{e}_{0}(0)=\overline{e}_{0}(2)=0$ and $e_{1}(2)=\overline{e}_{1}(2)=2$. For $k\geq n$, we always have $e_{k}(n)=0$.
  4. (d)
    If $n>2$ and $k\geq 2$, we have
    1. (i)
      $\overline{e}_{0}(n)=e_{0}(n-1)$ and $e_{0}(n)=\sum_{m\in P_{0,n}}\prod_{j=1}^{n}\binom{\overline{e}_{0}(j)+m_{0,j}-1}{m_{0,j}}$
    2. (ii)
      $\overline{e}_{1}(n)=e_{\Delta}(n)-\overline{e}_{0}(n)$ and $e_{1}(n)=\sum_{j=2}^{n}\overline{e}_{1}(j)e_{0}(n-j)$
    3. (iii)
      $\overline{e}_{k}(n)=\begin{cases}2,&\text{ if }k+1=n\text{ is even}\\ 1,&\text{ if }k+1=n\text{ is odd}\\ 0,&\text{ else.}\end{cases}$
    4. (iv)
[…]
Proof.
Recall that by Marshall’s normalized decomposition in 2.4, every quaternionic structure $S$ decomposes as $\mathbb{L}_{1,\varepsilon}^{k}\times S_{1}\times\ldots\times S_{\ell}$ for unique nondegenerate indecomposable quaternionic structures $S_{1},\ldots,S_{\ell}$ (up to permutation and isomorphism) and unique $k\in\mathbb{N}_{0}$. Further, we have $\varepsilon=0$ if and only if $s(S)=1$ and $\varepsilon=1$ else.
(a): The degenerate structures are clearly the ones having $k>0$ in the above notation. If $S$ is a degenerate structure of level $1$, it can uniquely be written as $S\cong\mathbb{L}_{1,0}\times S^{\prime}$ for $S^{\prime}\in\mathcal{E}^{\prime}_{n-1}$. This shows that there are $e^{\prime}(n-1)$ degenerate structures of level $1$ and order $n$. If $S$ is a degenerate structure of level different from $1$, we uniquely write $S\cong\mathbb{L}_{1,1}\times S^{\prime\prime}$ for a structure $S^{\prime\prime}$ of order $n-1$ that is either of level different from $1$ or both nondegenerate and of level $1$. Since the number of degenerate structures of level $1$ and order $n-1$ is $e^{\prime}(n-2)$ by the above argument, there are $(e(n-1)-e^{\prime}(n-1))+(e^{\prime}(n-1)-e^{\prime}(n-2))=e(n-1)-e^{\prime}(n-2)$ choices for $S^{\prime\prime}$.
(b): The only decomposable group extension is $\mathbb{L}_{1}\Delta\cong\mathbb{L}_{1}^{2}$, so the result follows from the fact that $S\Delta\cong S^{\prime}\Delta\iff S\cong S^{\prime}$ for any quaternionic structures $S,S^{\prime}$ of order greater than $2$.
(c): This is clear from the definitions and the list of small quaternionic structures (cf. Example 2.6).
(d) (d)(i): The set $\mathcal{E}_{0}(n)$ contains exactly the pythagorean quaternionic structures of order $n$. Since $n>2$ and local structures are not pythagorean, such a structure is indecomposable if and only if it is the group extension of a pythagorean structure of order $n-1$. Hence, $\overline{e}_{0}(n)=e_{0}(n-1)$. For the second part, we have to count the number of products of indecomposable pythagorean structures of smaller orders that are adding up to $n$. In other words, we just count the number of possible products for each partition $(m_{0,1},\ldots,m_{0,n})\in P_{0,n}$ of the order $n$, we can choose $m_{0,i}$ factors from $\overline{\mathcal{E}_{0}}(i)$ for all $i$. This gives the desired equation.
(d)(ii): The first part follows immediately as the strictly rigid quaternionic structures are exactly the group extensions that are not pythagorean (cf. results stated in Proposition 2.8). For the second part, note that the nondegenerate strictly rigid quaternionic structures are exactly those with one strictly rigid factor in the normalized decomposition while all other factors have to be pythagorean. Hence, such a structure can be uniquely written as the product of an indecomposable strictly rigid structure of order $j\in\{2,\ldots,n\}$ ($j\not=1$ as the factors are nondegenerate) and a pythagorean structure of order $n-j$.
(d)(iii): The indecomposable elementary type structures that have no rigid or 00-rigid elements are exactly the local structures. They have only two quaternions and are nondegenerate which immediately shows that local structures of order $n$ are strictly $(n-1)$-rigid.
(d)(iv): Clearly, for $S_{\ell}\in\mathcal{E}_{\ell}$ and $S_{\ell^{\prime}}\in\mathcal{E}_{\ell^{\prime}}$, we get $S_{\ell}\times S_{\ell^{\prime}}\in\mathcal{E}_{\ell+\ell^{\prime}}$. Hence, we have to count all products of nondegenerate indecomposable structures $S_{i,j}$ from $\mathcal{E}_{i}(j)$ such that $\sum_{i}i=k$ and $\sum_{j}j=n$. By definition of $P_{k,n}$, the desired equation follows in the same way as in (d)(i). ∎
Remark 2.10.
Lemma 2.9 enables us to refine the list Carson and Marshall gave in [CarsonMarshall_Decomposition] calculating $e(n)$ and $e^{\prime}(n)$ for small values of $n$, see Table 2. These values will be used later on for the computations in 6. We get the following:
$n$ 0 1 2 3 4 5 6 7 8 9
$e(n)$ 1 3 6 17 51 155 492 1600 5340 18150
$e^{\prime}(n)$ 1 1 2 4 10 22 54 130 328 832
$e_{\text{d}}(n)$ 0 2 3 7 19 57 167 524 1676 5538
$e_{0}(n)$ 1 1 1 2 4 8 17 36 79 175
$e_{1}(n)$ 0 0 2 7 22 71 227 735 2414 8059
$e_{2}(n)$ 0 0 0 1 4 14 60 234 903 3379
$e_{3}(n)$ 0 0 0 0 2 4 13 45 173 667
$e_{4}(n)$ 0 0 0 0 0 1 6 19 65 231
$e_{5}(n)$ 0 0 0 0 0 0 2 6 19 65
$e_{6}(n)$ 0 0 0 0 0 0 0 1 9 27
$e_{7}(n)$ 0 0 0 0 0 0 0 0 2 8
$e_{8}(n)$ 0 0 0 0 0 0 0 0 0 1
Table 2. Small values of $e(n),e^{\prime}(n),e_{d}(n),e_{k}(n)$
3. The Abstract 2-Brauer Group
To properly deal with quaternionic structures, we need some structure on the set of quaternions. For this, we imitate the $2$-torsion part of the Brauer group.
Definition.
Let $S=(G,-1,q)$ be a quaternionic structure. Let $(B,\ast)$ be a group into which the set of quaternions $Q$ embeds such that
(LR) $q(a,b)\ast q(a,c)=q(a,bc)\quad\text{for any }a,b,c\in G$
and $B$ is generated by $Q$. We then call $B$ a 2B-group of $S$. We call (LR) the linkage relation and say that two quaternions $q(a,b)$ and $q(c,d)$ are linked if there are $x,y,z\in G$ such that $q(a,b)=q(x,y)$ and $q(c,d)=q(x,z)$. The free group $\overline{Q}$ generated by the set $Q$ with the linkage relations as its sole relations trivially is a 2B-group and will be called the full 2B-group of $S$.
Lemma 3.1.
Let $S=(G,-1,q)$ be a quaternionic structure with 2B-group $B$ and set of quaternions $Q$. The following hold:
  1. (a)
    $B$ is a group of exponent $2$ with neutral element 00.
  2. (b)
    For any $\alpha\in\mathbb{N}_{0}$ and $a,b\in G$, we have $q(a^{\alpha},b)=q(a,b)^{\alpha}=q(a,b^{\alpha})$.
  3. (c)
    If $S$ is of order $n\in\mathbb{N}_{0}$, then
    $\dim(B)\leq\begin{cases}\binom{n}{2}&\text{ if }1=-1\\ \binom{n}{2}+1&\text{ else.}\end{cases}$
Proof.
(a): By definition, we have $q(a,b)\ast q(a,b)=q(a,b^{2})=q(a,1)=0$ and $q(a,b)\ast 0=q(a,b)\ast q(a,1)=q(a,b)$, so $B$ is a group of exponent $2$ with neutral element 00.
(b): Since both $G$ and $B$ are groups of exponent $2$, we can reduce to the case $\alpha\in\{0,1\}$. If $\alpha=0$, then all quaternions are trivial by Remark 1.1. If $\alpha=1$, all exponents are trivial, so there is nothing to show.
(c): Choose a basis $(a_{0},\ldots,a_{n-1})$. Then $B$ is generated by all $q(\prod_{i\in I}a_{i},\prod_{j\in J}a_{j})$ for subsets $I,J\subset\{0,\ldots,n-1\}$. Iteratively applying the linkage relation, we get
[…]
so $B$ is generated by the $q(a_{i},a_{j})$. By symmetry (Q2), we can always assume $i\leq j$ here. Furthermore, choosing $a_{0}=-1$ if $-1\not=1\in G$, for any $i\in\{0,\ldots,n-1\}$, we get (cf. Remark 1.1)
$q(a_{i},a_{i})=q(-1,a_{i})=\begin{cases}q(a_{0},a_{i})&\text{ if }-1\not=1\\ 0&\text{ else.}\end{cases}$
Hence, $B$ is generated by $(q(a_{i},a_{j})\mid 0\leq i<j\leq n-1)$ and, if $1$-1$\not$=1- 1 ≠ 1 additionally $q(a_{0},a_{0})$. The assertion follows. ∎
Proposition 3.2.
Let $S=(G,-1,q)$ and $S^{\prime}=(G^{\prime},-1^{\prime},q^{\prime})$ be quaternionic structures with full 2B-groups $B$ and $B^{\prime}$. The following hold:
  1. (a)
    The full 2B-group of $S\times S^{\prime}$ is isomorphic to $B\times B^{\prime}$.
  2. (b)
    The full 2B-group of $S\Delta$ is isomorphic to $B\times G$.
Proof.
(a): Let $Q$ respectively $Q^{\prime}$ be the sets of quaternions of $S$ and $S^{\prime}$. By Proposition 2.1, the set of quaternions of $S\times S^{\prime}$ can be chosen as $Q\times Q^{\prime}$. The linkage relations for $S\times S^{\prime}$ are thus exactly the linkage relations for $S$ and $S^{\prime}$ and it follows that the full 2B-group of $S\times S^{\prime}$ is isomorphic to $B\times B^{\prime}$.
(b): Write $S\Delta=S[x]=(G\cup Gx,-1,q_{x})$, $Q_{x}$ for its set of quaternions and $B_{x}$ for the full 2B-group. Take $a,b\in G$ and $\alpha,\beta\in\{0,1\}$. Again by Proposition 2.1, the elements of $Q_{x}$ can be viewed as a subset of $Q\times G$ with zero-element $(0,1)$, identifying $q_{x}(ax^{\alpha},bx^{\beta})$ with $(q(a,b),(-1)^{\alpha\beta}a^{\beta}b^{\alpha})$. Making use of Remark 1.1 and Lemma 3.1 (c) along with the linkage relations for $q_{x}$, we also get
$q_{x}(ax^{\alpha},bx^{\beta})$ $=q_{x}(a,b)\ast q_{x}(a,x^{\beta})\ast q_{x}(b,x^{\alpha})\ast q_{x}(x^{\alpha},x^{\beta})$
$=q_{x}(a,b)\ast q_{x}(a^{\beta}b^{\alpha},x)\ast q_{x}(x,x)^{\alpha\beta}$
$=q_{x}(a,b)\ast q_{x}(a^{\beta}b^{\alpha},x)\ast q_{x}(-1,x)^{\alpha\beta}$
$=q_{x}(a,b)\ast q_{x}((-1)^{\alpha\beta}a^{\beta}b^{\alpha},x).$
Hence, $B_{x}$ is generated by $\{q_{x}(a,b),q_{x}(a,x)\mid a,b\in G\}$. We define
$B^{\prime}:=\operatorname{span}(q_{x}(a,b)\mid a,b\in G)\text{ and }G^{\prime}:=\operatorname{span}(q_{x}(a,x)\mid a\in G).$
Since $q_{x}(a,x)\ast q_{x}(b,x)=q_{x}(ab,x)$ for all $a,b\in G$, we have $G^{\prime}\cong G$. It remains to show that $B^{\prime}\cong B$ and $B^{\prime}\cap G^{\prime}$ is trivial. For the former, take $a,b,c,d,e,f\in G$ and $\alpha,\beta\in\{0,1\}$. Since we only have the linkage relations, it follows that
(1) $\quad\quad q_{x}(a,b)\ast q_{x}(c,d)=q_{x}(ex^{\alpha},fx^{\beta})$
if and only if there exist $r,s,t\in G$ and $\rho,\sigma,\tau\in\{0,1\}$ such that
$q_{x}(a,b)=q_{x}(rx^{\rho},tx^{\tau}),\quad q_{x}(c,d)=q_{x}(sx^{\sigma},tx^{\tau}),\quad q_{x}(ex^{\alpha},fx^{\beta})=q_{x}(rsx^{\rho+\sigma},tx^{\tau}).$
Using the identification from Proposition 2.1, this is equivalent to the existence of such $r,s,t,\rho,\sigma,\gamma$ satisfying
$q(a,b)=q(r,t),\ \ q(c,d)=q(s,t),\ \ q(e,f)=q(rs,t)\text{ and}$
$(-1)^{\rho\tau}r^{\tau}t^{\rho}=1,\ \ (-1)^{\sigma\tau}s^{\tau}t^{\sigma}=1,\ \ (-1)^{(\rho+\sigma)\tau}(rs)^{\tau}t^{\rho+\sigma}=(-1)^{\alpha\beta}e^{\beta}f^{\alpha}.$
The term on the right hand side is the product of the two previous ones, so $(-1)^{\alpha\beta}e^{\beta}f^{\alpha}=1$. If $\alpha=1$, this means that $f=(-e)^{\beta}$, so either $\beta=0$ and $f=1$, meaning $q_{x}(ex^{\alpha},fx^{\beta})=q_{x}(ex,1)$, or $\beta=1$ and $q_{x}(ex^{\alpha},fx^{\beta})=q_{x}(ex,-ex)$. In both cases $q_{x}(ex^{\alpha},fx^{\beta})$ is the trivial quaternion $q_{x}(1,1)$. All in all, this shows that (1) holds if and only if
$q(a,b)\ast q(c,d)=q(e,f)\text{ and }(\alpha=\beta=0\text{ or }q_{x}(a,b)=q_{x}(c,d)).$
Since by Proposition 2.1, also $q_{x}(a,b)=q_{x}(c,d)\iff q(a,b)=q(c,d)$, it follows that $B^{\prime}$ is generated by the set $\{q_{x}(a,b)\mid a,b\in G\}$ with the exactly same (linkage) relations as $B$ with generating set $\{q(a,b)\mid a,b\in G\}$ and thus $B\cong B^{\prime}$. Additionally, the above calculations show that the only element of $G^{\prime}$ that is generated by $\{q_{x}(a,b)\mid a,b\in G\}$ is the trivial quaternion, so $B^{\prime}\cap G^{\prime}$ is trivial which remained to prove. ∎
Corollary 3.3.
Both upper bounds in Lemma 3.1(c) can be obtained, taking the iterated group extensions $\mathbb{L}_{0}\Delta_{n}$ or $\mathbb{L}_{1}\Delta_{n-1}$.
Proof.
Clearly, the full 2B-group of $\mathbb{L}_{0}$ has dimension 0, so it follows from Proposition 3.2 that the full 2B-group of $\mathbb{L}_{0}\Delta_{n}$ has dimension
$0+0+1+\ldots+(n-1)=\binom{n}{2}.$
The full 2B-group of $\mathbb{L}_{1}$ has dimension 1 as the unique nonzero quaternion necessarily must form a basis. Hence, the full 2B-group of $\mathbb{L}_{1}\Delta_{n-1}$ has dimension
$1+1+\ldots+(n-1)=\binom{n}{2}+1.\qed$
If $F$ is a field and $I$ the fundamental ideal of its Witt ring, the factor group $I^{2}/I^{3}$ is isomorphic to the $2$-torsion part of the Brauer group by Merkurjev’s Theorem [Merkurjev_KTheory]. In our abstract setting, the full $2B$-group will function as a substitute for $I^{2}/I^{3}$. We do so to avoid assuming $AP_{3}$ as an additional property of quaternionic structures. In the remainder of this section, we show that both $I^{2}/I^{3}$ and the full 2B-group are in fact the same for elementary types.
Proposition 3.4.
Let $S$ be an $AP_{3}$ structure with fundamental ideal $I$. Then $I^{2}/I^{3}$ is a 2B-group.
Proof.
As pointed out in Remark 1.3, the $k$-th power of the fundamental ideal is additively generated by the $k$-fold Pfister forms. In particular, the set of quaternions generates $I^{2}$, viewing $q(a,b)$ as $\langle\!\langle a,b\rangle\!\rangle=\langle 1,-a\rangle\langle 1,-b\rangle$ in the Witt ring $W$ of $S$ for $a,b\in G$. Assuming $AP_{3}$, we have to show that $Q$ also embeds into $I^{2}/I^{3}$. To do so, we show that
$q(a,b)=q(c,d)\in Q\iff\langle 1,-a\rangle\langle 1,-b\rangle-\langle 1,-c\rangle\langle 1,-d\rangle\in I^{3}\quad\text{for any }a,b,c,d\in G.$
The implication from left to right is trivial. If $\langle 1,-a\rangle\langle 1,-b\rangle-\langle 1,-c\rangle\langle 1,-d\rangle\in I^{3}$, then the Albert form $\varphi:=\langle-a,-b,ab,c,d,-cd\rangle$ lies in $I^{3}$. By $AP_{3}$, it follows that $\varphi$ must be hyperbolic, proving the claim.
It remains to prove the linkage relation: For any $a,b,c\in G$, we have
$\langle 1,-a\rangle\langle 1,-b\rangle+\langle 1,-a\rangle\langle 1,-c\rangle-\langle 1,-a\rangle\langle 1,-bc\rangle$ $=\langle 1,-a\rangle\langle 1,-b,-c,bc\rangle$
$=\langle 1,-a\rangle\langle 1,-b\rangle\langle 1,-c\rangle\in I^{3},$
completing the proof. ∎
Definition.
Let $S$ be an $AP_{3}$ structure with fundamental ideal $I$. We then call $I^{2}/I^{3}$ the (abstract) 2-Brauer group of $S$.
For the reader’s convenience, we recall Marshall’s results from [Marshall_AbstractWittRings, proof of 5.7, 5.17] on how the Arason-Pfister property $AP_{k}$ behaves for direct products and group extensions. Note that the statement for the direct product is only proven there for $k\in\{1,2\}$ but the proof works in the same way for larger $k$.
Proposition 3.5.
Let $S$ and $S^{\prime}$ be quaternionic structures and $k\geq 2$
  1. (a)
    If $S$ and $S^{\prime}$ both satisfy $AP_{k}$, then $S\times S^{\prime}$ also satisfies $AP_{k}$.
  2. (b)
    If $S$ satisfies both $AP_{k}$ and $AP_{k-1}$, then $S\Delta$ satisfies $AP_{k}$.
Proposition 3.6.
Let $S=(G,-1_{G},q)$ and $S^{\prime}=(G^{\prime},-1_{G^{\prime}},q^{\prime})$ be $AP_{3}$ structures with abstract Brauer groups $B$ and $B^{\prime}$. Then $S\times S^{\prime}$ and $S\Delta$ are $AP_{3}$ structures with abstract 2-Brauer groups
$B_{S\times S^{\prime}}\cong B\times B^{\prime}\quad\text{ and }\quad B_{S\Delta}\cong B\times G.$
Proof.
The fact that both $S\times S^{\prime}$ and $S\Delta$ are $AP_{3}$ structures follows directly from Proposition 3.5 along with the fact that $AP_{1}$ and $AP_{2}$ hold for any quaternionic structure (cf. Remark 1.3). Thus, $B_{S\times S^{\prime}}$ and $B_{S\Delta}$ are well-defined. We first consider the direct product. Write $W$ and $W^{\prime}$ for the Witt rings of $S$ and $S^{\prime}$. The Witt ring of $S\times S^{\prime}$ is the subring of $W\times W^{\prime}$ consisting of all $(\varphi,\psi)$ such that $\dim(\varphi)\equiv_{2}\dim(\psi)$. Since there are only even dimensional forms in the 2-Brauer groups, it immediately follows that the 2-Brauer group of $S\times S^{\prime}$ must be isomorphic to $B\times B^{\prime}$.
Now we move on to the group extension which we henceforth denote as $S[x]=(G\cup Gx,-1,q_{x})$. We write $I_{x}$ for its fundamental ideal. Note that we can write any $2$-fold Pfister form in $S[x]$ as $\langle\!\langle a,b\rangle\!\rangle$ or $\langle\!\langle a,bx\rangle\!\rangle$ for $a,b\in G$ as $\langle\!\langle ax,bx\rangle\!\rangle\cong\langle\!\langle-ab,bx\rangle\!\rangle$. With the same notation, we have $\langle\!\langle a,b\rangle\!\rangle+\langle\!\langle a,x\rangle\!\rangle=\langle\!\langle a,bx\rangle\!\rangle$ as elements of $I_{x}^{2}/I_{x}^{3}$, applying the linkage relation. Comparing with Proposition 3.2(b), $I_{x}^{2}/I_{x}^{3}$ is generated by $H_{1}:=\{\langle\!\langle a,b\rangle\!\rangle\mid a,b\in G\}$ and $H_{2}:=\{\langle\!\langle a,x\rangle\!\rangle\mid a\in G\}$. Again by the linkage relation, $\operatorname{span}(H_{2})\cong G$. Since $H_{1}$ embeds into the abstract 2-Brauer group of $S$, we clearly get $\operatorname{span}(H_{1})\cong B$ and $\operatorname{span}(H_{1})\cap\operatorname{span}(H_{2})=\{0\}$. As in Proposition 3.2(b), the assertion follows. ∎
Corollary 3.7.
Let $S$ be an elementary type structure with set of quaternions $Q$ and 2-Brauer group $B$. Then $B\cong\overline{Q}$.
Proof.
By Proposition 3.2 and Proposition 3.6, it suffices to show this if $|Q|\leq 2$. If $|Q|=1$, then any group of exponent $2$ generated by $Q$ is trivial, so the assertion follows. If $|Q|=2$, there is also only one group of exponent $2$ generated by $Q$ because $Q$ is multiplicatively closed ($0\ast 0=0=1\ast 1$, $1\ast 0=1=0\ast 1$), so $B=Q=\overline{Q}$. ∎
4. Quaternionic Matrices
In the following, we only consider structures $S$ of finite order $n\in\mathbb{N}$. In this section, we establish a unique matrix $M_{S}\in\operatorname{Mat}_{n}(\mathbb{N}_{0})$ by which we can identify $S$. This will be the groundwork for our computations later on. We usually write $S=(G,-1,q)$ with set of quaternions $Q$ and full $2B$-group $B$. We call its dimension $\dim(S):=\dim(B)$, which is finite by Lemma 3.1(c), the dimension of $S$.
Notation.
For $n\in\mathbb{N}$, we often consider matrices $M=(M_{i,j})_{0\leq i,j\leq n-1}\in\operatorname{Mat}_{n}(\mathbb{N}_{0})$ where we start counting the indices for rows and columns at 00. We bestow the set $\operatorname{Mat}_{n}(\mathbb{N}_{0})$ with the lexicographical order $<_{\text{lex}}$, applied row by row, and extending the usual order on $\mathbb{N}_{0}$. We simply say that a matrix is ’smaller/greater/…’ than another if it is lexicographically smaller/greater/…
For $k\in\mathbb{N}_{0}$, we take $I_{k}\subset\mathbb{N}_{0}$ such that $k=\sum_{i\in I_{k}}2^{i}$ is the binary representation of $k$. Then the binary operation $\ast_{\mathbb{N}_{0}}$ on $\mathbb{N}_{0}$ where
$k\ast_{\mathbb{N}_{0}}l:=\sum_{i\in I_{k}\triangle I_{l}}2^{i}$
defines a group of exponent $2$ $(\mathbb{N}_{0},\ast_{\mathbb{N}_{0}})$ where any set $\{0,1,\ldots,2^{n}-1\}$ defines a subgroup with basis $(2^{k}\mid 0\leq k\leq n-1)$. This is easily verified using linear algebra. We often simply write $\ast$ instead of $\ast_{\mathbb{N}_{0}}$ and define the completion of $M$ as the matrix
[…]
Note that we usually write $\overline{M}_{I_{k},I_{\ell}}:=\overline{M}_{k,\ell}$, using subsets of $\{0,\ldots,n-1\}$ as indices. For $k=0$ or $\ell=0$, the above definition gives an empty product, meaning that the leading row and column of $\overline{M}$ consist only of zeros.
We also define $(B(M),\overline{\ast})$ as the free group of exponent $2$ $\langle\overline{M}_{I,J}\mid I,J\subset\{0,\ldots,n-1\}\rangle$ with the (linkage) relations
$\overline{M}_{I,J}\overline{\ast}\overline{M}_{I,K}=\overline{M}_{I,J\triangle K}$
for all $I,J,K\subset\{0,\ldots,n-1\}$.
Definition.
We say that a basis $\mathcal{A}=(a_{0},\ldots,a_{n-1})$ of $G$ is a basis of S if either $-1=1$ or $a_{0}=-1$. Given a basis $\mathcal{B}=(q_{0},\ldots,q_{m-1})$ of $B$, we define the group isomorphism
$\phi_{\mathcal{B}}:(B,\ast)\to(\{0,\ldots,2^{m}-1\},\ast)\quad\text{with}\quad q_{k}\mapsto 2^{k}.$
Then, we define the quaternionic matrix of $S$ via $\mathcal{A},\mathcal{B}$ as
$M_{\mathcal{A},\mathcal{B}}(S):=(\phi_{\mathcal{B}}(q(a_{i},a_{j})))_{0\leq i,j\leq n-1}\in\operatorname{Mat}_{n}(\mathbb{N}_{0}).$
If $\mathcal{B}$ is fixed, we often identify $(B,\ast)$ with ($(\{0,\ldots,2^{m}-1\},\ast)$, writing $q_{\mathcal{B}}(a,b)$ or – in abuse of notation – simply $q(a,b)$ instead of $\phi_{\mathcal{B}}(q(a,b))$.
Example 4.1.
Note that the matrices defined as above are in general not the incidence matrices as studied by Kula (cf. [Kula_FinitelyGeneratedWittRings]). The explicit matrices given here are in fact the lexicographically minimal ones for the respective structures which we leave for the reader to check. The minimality will be of importance for defining a unique quaternionic matrix of $S$.
  1. (a)
    If $|Q|=1$ (i.e. $S$ is totally degenerate), then $M_{\mathcal{A},\mathcal{B}}$ is the $n\times n$ zero matrix, independent of the choice of the basis $\mathcal{A}$ of $G$ with $\mathcal{B}$ being empty since $\dim(S)=0$.
  2. (b)
    If $|Q|=2$, there is again only one choice for $\mathcal{B}$ and for any basis $\mathcal{A}$ of $G$, the quaternionic matrix $M_{\mathcal{A},\mathcal{B}}$ is the representation matrix of the symmetric bilinear map $G\times G\to B$, identifying $B$ with $\mathbb{F}_{2}$. If $S$ is of local type, it is a linear algebra exercise to show that we can choose $\mathcal{A}$ in a way that
    • if $S\cong\mathbb{L}_{2k,0}$: $M_{\mathcal{A},\mathcal{B}}$ is the matrix $E_{2k}\in M_{2k}(\mathbb{N}_{0})$ with entry $1$ on the anti-diagonal and entry 00 else.
    • if $S\cong\mathbb{L}_{2k,1}$:
      $M_{\mathcal{A},\mathcal{B}}=\begin{pmatrix}0&0&\ldots&0&1\\ 0&&&&0\\ \vdots&&E_{2k-2}&&\vdots\\ 0&&&&0\\ 1&0&\ldots&0&1\end{pmatrix}$
    • if $S\cong\mathbb{L}_{2k-1}$:
      $M_{\mathcal{A},\mathcal{B}}=\begin{pmatrix}1&0&\ldots&0\\ 0&&&\\ \vdots&&E_{2k-2}&\\ 0&&&\end{pmatrix}$
    • We include a list of quaternionic matrices for the nondegenerate quaternionic structures of order $2$ which can be computed similarly to the previous example. We do not explicitly state the choice of bases as it can be checked that the matrices are mostly independent of that choice with the exception of $\mathbb{L}_{1}\Delta$, where we have some liberties regarding the choice of $\mathcal{B}$ which however does not change the reduced version of the matrix (cf. later on in Lemma 4.11).
      $\mathbb{L}_{1,0}\Delta:\begin{pmatrix}0&1\\ 1&0\end{pmatrix},\quad\mathbb{L}_{1,1}\Delta:\begin{pmatrix}0&1\\ 1&1\end{pmatrix},\quad\mathbb{L}_{1}\Delta:\begin{pmatrix}1&2\\ 2&2\end{pmatrix},$
Lemma 4.2.
Let $\mathcal{A}=(a_{0},\ldots,a_{n-1})$ be a basis of $S$ and $\mathcal{B}$ a basis of $B$. Set $M:=M_{\mathcal{A},\mathcal{B}}(S)$. The following hold:
  1. (a)
    For any $I,J\subset\{0,\ldots,n-1\}$, we have $\overline{M}_{I,J}=q_{\mathcal{B}}(\prod_{i\in I}a_{i},\prod_{j\in J}a_{j})$.
  2. (b)
    $M$ satisfies the following:
    1. (M1)
      $B(M)\cong\operatorname{span}_{\ast_{\mathbb{N}_{0}}}(M_{i,j}\mid 0\leq i,j\leq n-1)$
    2. (M2)
      $M$ is symmetric
    3. (M3)
      $M_{i,i}=M_{0,i}$ for all $i\in\{0,\ldots,n-1\}$ or $M_{i,i}=0$ for all $i\in\{0,\ldots,n-1\}$
    4. (M4)
      For all $I,J,K,L\subset\{0,\ldots,n-1\}$:
      If $\overline{M}_{I,J}=\overline{M}_{K,L}$, there is $X\!\subset\!\{0,\ldots,n-1\}$ with $\overline{M}_{I,J}=\overline{M}_{I,X}=\overline{M}_{K,X}=\overline{M}_{K,L}.$
  3. (c)
    The basis $\mathcal{B}$ can be chosen such that $M$ also satisfies
    1. (M0)
      For any $i,j\in\{0,\ldots,n-1\}$: $M_{i,j}\leq\operatorname{min}\{2^{k}\mid M_{i^{\prime},j^{\prime}}<2^{k}\ \forall\ (i^{\prime},j^{\prime})<_{\text{lex}}(i,j)\})$.
    In this case, we have $M_{\mathcal{A},\mathcal{B}}(S)\leq_{\text{lex}}M_{\mathcal{A},\mathcal{B}^{\prime}}$ for all bases $\mathcal{B}^{\prime}$ of $B$ and define $M_{\mathcal{A}}(S):=M$, the reduced quaternionic matrix of $S$ via $\mathcal{A}$.
Proof.
Write $m:=\dim(S)$.
(a): Since $\phi_{\mathcal{B}}$ is an isomorphism, we have
[…]
and since $q_{\mathcal{B}}(a_{i},a_{j})=M_{i,j}$, the above equals $\overline{M}_{I,J}$ by definition.
(b): The properties (b)(M2) and (b)(M4) follow directly from (Q2) and (Q4). For (b)(M3), it suffices to see that $q(a,a)=q(a,-1)$ for all $a\in G$ by Remark 1.1 and this equals 00 if $1=-1$ in $G$. Since we have chosen $a_{0}=-1$ in the other case, (b)(M3) follows.
For (b)(M1), recall that the full $2B$-group $B$ of $S$ is the free group of exponent $2$ generated by $Q$ with the linkage relations as its sole relations. The linkage relation directly translates to the relation given in the definition of $B(M)$ by (a). As $B$ is generated by the $q(a_{i},a_{j})$, we get the desired property.
(c): Since the $q(a_{i},a_{j})$ generate $B$ (cf. Lemma 3.1), we can choose $\mathcal{B}$ from these quaternions. Take $k\in\mathbb{N}_{0}$. If $q_{0},\ldots,q_{k-1}$ are defined and $k<m$, set $q_{k}:=q(a_{i_{k}},a_{j_{k}})$ such that $(i_{k},j_{k})$ is lexicographically minimal such that $q(a_{i_{k}},a_{j_{k}})\not\in\operatorname{span}(q_{0},\ldots,q_{k-1})$. This iteratively defines a basis $\mathcal{B}=(q_{0},\ldots,q_{m-1})$ of $B$. To show (c)(M0), take $i,j\in\{0,\ldots,n-1\}$. If $(i,j)<_{\text{lex}}(i_{0},j_{0})$, then $q(a_{i},a_{j})=0$ and thus also $M_{i,j}=0$. If $(i,j)>_{\text{lex}}(i_{m-1},j_{m-1})$, there is nothing to show because all entries of $M$ are in $\{0,\ldots,2^{m}-1\}=\operatorname{span}(q_{0},\ldots,q_{m-1})$.
Otherwise, we can take $k\in\{0,\ldots,m-1\}$ minimal such that $(i_{k},j_{k})\geq_{\text{lex}}(i,j)$. Then for $(i^{\prime},j^{\prime})<_{\text{lex}}(i,j)$, we have $q(a_{i^{\prime}},a_{j^{\prime}})\in\operatorname{span}(q_{0},\ldots,q_{k-1})$ by construction of $\mathcal{B}$ and thus
$M_{i^{\prime},j^{\prime}}=q_{\mathcal{B}}(a_{i^{\prime}},a_{j^{\prime}})\in\operatorname{span}(1,2,\ldots,2^{k-1})=\{0,\ldots,2^{k}-1\}.$
If $(i,j)=(i_{k},j_{k})$, then $M_{i,j}=2^{k}$. Otherwise it follows as for $(i^{\prime},j^{\prime})$ above that $M_{i,j}<2^{k}$. If $k\neq 0$, then $2^{k-1}=M_{i_{k-1},j_{k-1}}$ and $i_{k-1},j_{k-1}<(i,j)$, so $k$ is the minimum as described in (c)(M0).
It follows from the construction that this choice of $\mathcal{B}$ gives the lexicographical minimal $M_{\mathcal{A},\mathcal{B}}(S)$ for fixed $\mathcal{A}$. ∎
Definition.
Take $n\in\mathbb{N}$ and $M\in\operatorname{Mat}_{n}(\mathbb{N}_{0})$. We call $M$ a prequaternionic matrix if it satisfies (b)(M2), (b)(M3), (b)(M4). If $M$ additionally satisfies (b)(M1), we say that $M$ is a quaternionic matrix and if it additionally satisfies (c)(M0) and (b)(M1), a reduced quaternionic matrix. If $M$ is a (pre-)quaternionic matrix, we say that $M$ is of type 1 if the first condition of (b)(M3) is satisfied and of type 0 if the second condition of (b)(M3) is satisfied.
Lemma 4.2 shows that in the notation used there, $M_{\mathcal{A},\mathcal{B}}(S)$ is a quaternionic matrix and $M_{\mathcal{A}}(S)$ is a reduced quaternionic matrix.
In the following, we need some technicalities to assure getting a unique representation of $S$ by a quaternionic matrix. First, we establish how prequaternionic matrices already uniquely define quaternionic structures up to isomorphism. Then, we use Remark 4.6 to illustrate why we nonetheless need (b)(M1) to properly reflect the structure of the full 2B-group.
Lemma 4.3.
Let $M\in\operatorname{Mat}_{n}(\mathbb{N}_{0})$ be a prequaternionic matrix of type $\epsilon\in\{0,1\}$ with completion $\overline{M}$. Then the following hold:
  1. (a)
    $\overline{M}$ is symmetric.
  2. (b)
    If $\epsilon=0$, then $\overline{M}_{I,I}=0$ for all $I\subset\{0,\ldots,n-1\}$.
    If $\epsilon=1$, then $\overline{M}_{I,I}=\overline{M}_{\{0\},I}$ for all $I\subset\{0,\ldots,n-1\}$.
  3. (c)
    $\overline{M}_{I,J}\ast\overline{M}_{I,K}=\overline{M}_{I,J\triangle K}$ for all $I,J,K\subset\{0,\ldots,n-1\}$.
  4. (d)
    $\overline{M}_{\{i\},\{j\}}=M_{i,j}$ for all $i,j\in\{0,\ldots,n-1\}$.
Proof.
(a) is an immediate consequence of (b)(M2) and the definition of $\overline{M}$. Now take $I,J,K\subset\{0,\ldots,n-1\}$. We get (b) by
[…]
In a similar fashion, we get (c) by computing
[…]
Lastly, (d) follows directly from the definition. ∎
The following shows that any prequaternionic matrix defines a quaternionic structure from which we can then recover the zeros in the completion of the matrix.
Proposition 4.4.
Let $M\in\operatorname{Mat}_{n}(\mathbb{N}_{0})$ be a prequaternionic matrix of type $\epsilon\in\{0,1\}$. Let $G$ be a group of exponent $2$ with neutral element $1$ and basis $\mathcal{A}:=(a_{0},\ldots,a_{n-1})$. If $\epsilon=0$, choose $-1:=1\in G$, else $-1:=a_{0}$. Define $Q$ as the set of entries of $\overline{M}$, naturally containing 0. Then
$q:G\times G\to Q,\quad\left(\prod_{i\in I}a_{i},\prod_{j\in J}a_{j}\right)\mapsto\overline{M}_{I,J}$
defines a quaternionic structure $S(M)=(G,-1,q)$ with $s(S)=1$ if $\epsilon=0$ and $s(S)>1$ else. For any basis $\mathcal{B}$ of the full $2B$-group of $S(M)$ and $I,J\subset\{0,\ldots,n-1\}$, it satisfies
$\overline{M}_{I,J}=0\iff\overline{M_{\mathcal{A},\mathcal{B}}(S(M))}_{I,J}=0$
Proof.
The property (Q2) follows from Lemma 4.3(a). (Q1) follows from Lemma 4.3(b), in the case $\epsilon=1$ also making use of part (c):
$q\left(\prod_{i\in I}a_{i},-\prod_{i\in I}a_{i}\right)=\overline{M}_{I,I\triangle\{0\}}=\overline{M}_{I,I}\ast\overline{M}_{I,\{0\}}=0.$
(Q3) is another consequence of Lemma 4.3(c), applying the fact that $(\mathbb{N}_{0},\ast)$ is a group of exponent $2$. Lastly, (Q4) follows from (b)(M4).
The zeros in the completion of a quaternionic matrix, for fixed basis $\mathcal{A}$, clearly do not depend on the choice of the basis of the full $2B$-group. ∎
Corollary 4.5.
  1. (a)
    Let $M,M^{\prime}\in\operatorname{Mat}_{n}(\mathbb{N}_{0})$ be prequaternionic matrices of the same type such that $\overline{M}$ and $\overline{M^{\prime}}$ have zeros in the same entries. Then $S(M)\cong S(M^{\prime})$ as quaternionic structures.
  2. (b)
    Let $S$ be a finite quaternionic structure, $\mathcal{A}$ a basis of $S$ and $\mathcal{B}$ a basis of the full $2B$-group of $S$. Then $S\cong S(M_{\mathcal{A},\mathcal{B}}(S))$.
Proof.
(a): Write $\mathcal{A}$ and $\mathcal{A}^{\prime}$ for the bases of $S$ and $S^{\prime}$ as in Proposition 4.4. It is clear that the basis transformation $\mathcal{A}\mapsto\mathcal{A}^{\prime}$ defines an isomorphism of quaternionic structures as it preserves the $(-1)$-element ($M,M^{\prime}$ are of the same type) and the zero quaternions.
(b): As before, this follows from the fact that the canonical isomorphism of groups of exponent $2$ preserves $-1$ and the zero quaternions. ∎
Remark 4.6.
At this point, we could simply assign a unique prequaternionic matrix to any quaternionic structure $S$ by choosing the lexicographically minimal one giving rise to a structure isomorphic to $S$ by the construction of Lemma 4.3. This would give a bijective mapping from the isomorphy types of quaternionic structures of order $n$ to a certain subset of $\operatorname{Mat}_{n}(\mathbb{N}_{0})$. As it turns out, this does not properly display the structure of the full $2B$-group of $S$: In general, even if additionally requiring (c)(M0), there can be prequaternionic matrices of $S$ that cannot be written as the (reduced) quaternionic matrix of $S$ via a basis of $S$. This is the case if there are relations of $\ast_{\mathbb{N}_{0}}$ appearing in the prequaternionic matrix that are not induced by linkage relations. We illustrate this in the following Example 4.7.
Note that for the matrices given in Example 4.1, it can be verified that (c)(M0) and (b)(M1) both hold, so these are in fact reduced quaternionic matrices.
Example 4.7.
To make notation a bit clearer in this example, we use some coloring to point out which elements belong to which ’part’ of the following structure: We consider the elementary type structure
$(\mathbb{L}_{1}\times({\color[rgb]{0,1,0}\mathbb{L}_{1}}\times{\color[rgb]{0,0,1}\mathbb{L}_{1}}){\color[rgb]{1,0,0}\Delta}){\color[rgb]{1,.5,0}\Delta}=:S=:(G,-1,q)$
with square class group $G=\{1,a_{1}\}\times{\color[rgb]{0,1,0}\{1,b_{0}\}}\times{\color[rgb]{0,0,1}\{1,a_{2}\}}\times{\color[rgb]{1,0,0}\{1,a_{3}\}}\times{\color[rgb]{1,.5,0}\{1,a_{4}\}}$ according to construction and coloring. Then $a_{0}:=-1=a_{1}{\color[rgb]{0,1,0}b_{0}}{\color[rgb]{0,0,1}a_{2}}$. We compute $M:=M_{\mathcal{A}}(S)$ for the basis $\mathcal{A}=(a_{0},\ldots,a_{4})$. We have $q(a_{0},a)\not=0$ for all $a\in G\setminus\{1\}$ ($S$ is pythagorean), so all entries of the first row of $M$ must be independent, meaning that by (c)(M0), we have (apply Proposition 5.1 from the following section or a straightforward computation) the first row (1 2 4 8 16)124816(1\ 2\ 4\ 8\ 16)( 1 2 4 8 16 ). Furthermore, by the properties of the direct product $\mathbb{L}_{1}\times({\color[rgb]{0,1,0}\mathbb{L}_{1}}\times{\color[rgb]{0,0,1}\mathbb{L}_{1}}){\color[rgb]{1,0,0}\Delta}$, we get $q(a_{1},{\color[rgb]{0,0,1}a_{2}})=0=q(a_{1},{\color[rgb]{1,0,0}a_{3}})$. Applying (b)(M2) and (b)(M3) there are $4$ more values to compute:
$M=\begin{pmatrix}1&2&4&8&16\\ 2&2&0&0&x_{1}\\ 4&0&4&x_{2}&x_{3}\\ 8&0&x_{2}&8&x_{4}\\ 16&x_{1}&x_{3}&x_{4}&16\end{pmatrix}$
Iteratively applying Proposition 3.2 for the full $2B$-group, it follows that
$\dim(S)=1+(({\color[rgb]{0,1,0}1}+{\color[rgb]{0,0,1}1})+{\color[rgb]{1,0,0}2})+{\color[rgb]{1,.5,0}4}=9.$
Hence, we must have $x_{1}=2^{5}$, $x_{2}=2^{6}$, $x_{3}=2^{7}$, $x_{4}=2^{8}$ and thus
$M=\begin{pmatrix}1&2&4&8&16\\ 2&2&0&0&32\\ 4&0&4&64&128\\ 8&0&64&8&256\\ 16&32&128&256&16\end{pmatrix}.$
It is clear from the uniqueness of the first row and the position of the zero entries that this is the lexicographically minimal quaternionic matrix of $S$ via a basis of $S$. However, it can be verified with an elementary yet lengthy computation that the matrix
$M^{\prime}:=\begin{pmatrix}1&2&4&8&16\\ 2&2&0&0&32\\ 4&0&4&33&64\\ 8&0&33&8&128\\ 16&32&64&128&16\end{pmatrix}$
also satisfies (c)(M0), $M^{\prime}<_{\text{lex}}M$ and we have $\overline{M}_{I,J}=0\iff\overline{M^{\prime}}_{I,J}=0$ for all $I,J\subset\{0,\ldots,4\}$. Hence, by Corollary 4.5, we have $S\cong S(M^{\prime})$.
Here, the ’problem’ lies in the fact that $33=32\ast 1$ is no relation of the full $2B$-group of $S$ where the according values are independent, as $32\ast 1=33$ is not induced by linkage relations.
This illustrates why we need the seemingly technical condition (b)(M1) to assure that our quaternionic matrices properly reflect the properties of the full $2B$-group.
As illustrated by the previous example, the axiom (b)(M1) indeed assures that the properties of the full 2B-group are reflected by the quaternionic matrices:
Corollary 4.8.
If $M\in\operatorname{Mat}_{n}(\mathbb{N}_{0})$ is a reduced quaternionic matrix and $S(M)$ is defined as in Proposition 4.4, then the different $2$-powers appearing as entries of $M$ form a basis of the full $2B$-group of $S(M)$. In particular, if $M$ has a nonzero entry, then
$\dim(S(M))=\operatorname{max}\{k+1\mid\exists\ i,j\in\{0,\ldots,n-1\}:\ M_{i,j}=2^{k}\}.$
Proof.
By (c)(M0), we have $m\in\mathbb{N}$ such that $\mathcal{B}:=(1,2,\ldots,2^{m-1})$ consists of entries of $M$ and all entries of $\overline{M}$ are in $\{0,\ldots,2^{m}-1\}$. Hence, $\mathcal{B}$ is a basis of $\operatorname{span}_{\ast}(M_{i,j}\mid 0\leq i,j\leq n-1)$. Since $\mathcal{B}$ is independent in $(\mathbb{N}_{0},\ast)$ and all relations by which we defined $B(M)$ hold in $(\mathbb{N}_{0},\ast)$, $\mathcal{B}$ stays independent viewed as elements in $B(M)$. Now by (b)(M1), $\mathcal{B}$ must be a basis of $B(M)$. It is the group of exponent $2$ generated by the quaternions with the linkage relations as its sole relations, which is exactly how the full $2B$-group is defined. ∎
Definition.
We call two quaternionic matrices of the same type isomorphic if they define isomorphic quaternionic structures in the sense of Proposition 4.4. Let $M$ be a quaternionic matrix of type $\epsilon\in\{0,1\}$. We say that a quaternionic matrix $M$ is a normal quaternionic matrix if for every quaternionic matrix $M^{\prime}$ isomorphic to $M$, we have $M\leq_{\text{lex}}M^{\prime}$.
Also, we call $M$ nondegenerate if $S(M)$ is nondegenerate.
By this definition, we immediately get the following:
Corollary 4.9.
Let $\epsilon\in\{0,1\}$. For any $n\in\mathbb{N}$, we have a bijection between the set of normal quaternionic matrices of type $\epsilon$ in $\operatorname{Mat}_{n}(\mathbb{N}_{0})$ and the isomorphy types of quaternionic structures $S$ of order $n$ with $-1=1$ (if $\epsilon=0$) or $1$-1$\not$=1- 1 ≠ 1 (if $\epsilon=1$), mapping $M$ to $S(M)$.
Restricting ourselves to nondegenerate quaternionic structures, we can handle both types simultaneously by (b)(M3).
Corollary 4.10.
For any $n\in\mathbb{N}$, we have a bijection between the set of nondegenerate normal quaternionic matrices in $\operatorname{Mat}(\mathbb{N}_{0})$ and the isomorphy types of nondegenerate quaternionic structures of order $n$.
Proof.
Clearly, if $M$ is a nondegenerate quaternionic matrix, then the first row of $M$ has a nonzero entry. Thus, $M$ cannot satisfy both conditions in (b)(M3) simultaneously, and the type of $M$ is uniquely determined. ∎
In the following, we always write $M_{S}$ for the normal quaternionic matrix of a quaternionic structure $S$. This notation is justified by Corollary 4.9.
Definition.
Let $M\in\operatorname{Mat}_{n}(\mathbb{N}_{0})$ be a quaternionic matrix of type $\epsilon\in\{0,1\}$. We say that $M^{\prime}\in\operatorname{Mat}_{n}(\mathbb{N}_{0})$ is a simple transformation of $M$ if there are $i,j\in\{0,\ldots,n-1\}$, $i\not=j$ such that
  • We have $\epsilon=0$ or $i,j\not=0$ and:
    $M^{\prime}$ is obtained from $M$ by swapping the $i$-th row of $M$ with the $j$-th row and the $i$-th column with the $j$-th column
  • We have $\epsilon=0$ or $i\not=0$ and:
    $M^{\prime}_{i,k}=M_{i,k}\ast M_{j,k}=M^{\prime}_{k,i}$ for all $k\not=i$, $M^{\prime}_{i,i}=M_{i,i}\ast M_{j,j}$ and $M^{\prime}_{k,\ell}=M_{k,\ell}$ for all $k,\ell\in\{0,\ldots,n-1\}\setminus\{i\}$
The simple transformations above will briefly be denoted as $(i\leftrightarrow j)$ and $(i\mapsto i\ast j)$ respectively. We say that $M^{\prime}$ is a transformation of $M$ if it can be obtained from $M$ using a finite amount of simple transformations.
The simple transformations display basis transformations of quaternionic structures in the following way: Take a quaternionic structure $S=(G,-1,q)$ with basis $\mathcal{A}=(a_{0},\ldots,a_{n-1})$ and basis $\mathcal{B}$ of the full $2B$-group. As $G$ is an $\mathbb{F}_{2}$ vector space, we cannot scale $\mathcal{A}$ nontrivially, so if $\mathcal{A}^{\prime}$ is another basis of $S$, we can obtain it from $\mathcal{A}$ by an iteration of swapping elements $a_{i}\leftrightarrow a_{j}$ or replacing $a_{i}$ by $a_{i}a_{j}$. The following lemma shows that the matrix transformations defined above translate the effects of these basis transformations to the quaternionic matrices, going from $M_{\mathcal{A},\mathcal{B}}$ to $M_{\mathcal{A}^{\prime},\mathcal{B}}$ by an iteration of simple transformations. For the canonical choice of $\mathcal{B}$ (cf. Lemma 4.2(c)) in this setting, we have to consider reduced quaternionic matrices:
Lemma 4.11.
Let $M,M^{\prime}\in\operatorname{Mat}_{n}(\mathbb{N}_{0})$ be quaternionic matrices of the same type. The following hold:
  1. (a)
    There is a unique reduced quaternionic matrix $M_{\text{red}}$ such that:
    $\forall I,J\subset\{0,\ldots,n-1\}:\ \overline{M}_{I,J}=0\iff\overline{M_{\text{red}}}_{I,J}=0$
  2. (b)
    If $M$ is a normal quaternionic matrix, then $M$ is a reduced quaternionic matrix.
  3. (c)
    $M$ and $M^{\prime}$ are isomorphic if and only if there is a transformation $M^{\prime\prime}$ of $M$ such that $M^{\prime\prime}_{\text{red}}=M^{\prime}_{\text{red}}$.
Proof.
The entries of $M$ span the subspace $B:=B(M)$ of $(\mathbb{N}_{0},\ast)$ by (b)(M1). Set $m:=\dim(B)$.
(a): We can choose a basis $\mathcal{B}:=(M_{i_{0},j_{0}},\ldots,M_{i_{m-1},j_{m-1}})$ of $B$ in the same way as in Lemma 4.2(c). Then, the group isomorphism $B\to\{0,\ldots,2^{m}-1\},\ \mathcal{B}\mapsto(1,2,\ldots,2^{m-1})$ applied on the entries of $M$ defines a matrix $M_{\text{red}}$. Clearly, for any $I,J,K,L\subset\{0,\ldots,n-1\}$, we have
$\overline{M}_{I,J}=\overline{M}_{K,L}\iff\overline{M_{\text{red}}}_{I,J}=\overline{M_{\text{red}}}_{K,L}.$
This immediately shows that $M_{\text{red}}$ is a prequaternionic matrix. The axiom (c)(M0) follows as in the proof of Lemma 4.2(c). (b)(M1) is clear as $B(M)\cong\operatorname{span}_{\ast}(1,2,\ldots,2^{m-1})$ is preserved by the basis transformation we applied on $\mathcal{B}$. It follows that $M_{\text{red}}$ is a reduced quaternionic matrix with the desired property. The uniqueness follows in the same way as in Lemma 4.2(c) again.
(b): This follows immediately from (a) since by Corollary 4.5(a), we get $S(M)\cong S(M_{\text{red}})$ and clearly $M_{\text{red}}\leq_{\text{lex}}M$ (cf. Lemma 4.2(c).
(c): First, we point out that applying simple transformations $(i\leftrightarrow j)$ or $(i\mapsto i\ast j)$ just means swapping the $i$-th and $j$-th basis element in the induced basis of $S(M)$ or replacing $a_{i}$ by $a_{i}a_{j}$. Both are basis transformations of $S(M)$ as by the choice of $i,j$, the element $-1$ is fixed. In particular, simple transformations and by iteration also general transformations of $M$ are again quaternionic matrices of the same type that are isomorphic to $M$. By (a), reduction of quaternionic matrices also does not change the isomorphy type, so the ’if’ part of the statement holds.
Conversely, assume that $S:=S(M)=(G,-1,q)\cong S(M^{\prime})=(G^{\prime},-1^{\prime},q^{\prime})=:S^{\prime}$ with bases $\mathcal{A}=(a_{0},\ldots,a_{n-1})$ and $\mathcal{A}^{\prime}$ as in Proposition 4.4. Consider the isomorphism $\phi:G\to G^{\prime}$ of $S$ and $S^{\prime}$ and the preimages $a_{i}^{\prime\prime}:=\phi^{-1}(a_{i}^{\prime})$ for $i=0,\ldots,n-1$. They form a basis $\mathcal{A}^{\prime\prime}:=(a_{0}^{\prime\prime},\ldots,a_{n-1}^{\prime\prime})$ of $S$. Hence, we can obtain $\mathcal{A}^{\prime\prime}$ from $\mathcal{A}$ by a finite amount of elementary basis transformations. Write $M:=M_{\mathcal{A},\mathcal{B}}(S)$ for a basis $\mathcal{B}$ of the full $2B$-group of $S$. Then $M^{\prime\prime}:=M_{\mathcal{A}^{\prime\prime},\mathcal{B}}$ is a transformation of $M$. Since $\phi$ is an isomorphism mapping $\mathcal{A}^{\prime\prime}$ to $\mathcal{A}^{\prime}$, the completions $\overline{M^{\prime\prime}}$ and $\overline{M^{\prime}}$ have zeros at the same entries as in 4.4. By (a), this shows that $M^{\prime\prime}_{\text{red}}=M^{\prime}_{\text{red}}$. ∎
The previous statement points out that in order to compute the normal quaternionic matrix of a given quaternionic matrix $M$, we have to compute the lexicographic minimum of all reduced (in the sense of Lemma 4.11(a)) transformations of $M$.
5. Normal Quaternionic Matrices
The main result of this section will be determining what the first two rows of normal quaternionic matrices look like. The structure of the first row is determined quite easily:
Proposition 5.1.
Let $M=(M_{i,j})_{0\leq i,j\leq n-1}$ be a normal quaternionic matrix of type $\epsilon\in\{0,1\}$ with quaternionic structure $S:=S(M):=(G,-1,q)$. Set $x:=\dim(V(-1))$ if $-1\not=1\in G$ and otherwise $x:=\operatorname{max}\{\dim(V(a))\mid a\in G\setminus\{1\}\}$. We write $s:=s(S)$. Then
$(M_{0,0},M_{0,1},\ldots,M_{0,n-1})=\begin{cases}(0,\ldots,0,1,2,4,\ldots,2^{n-x-1}),&\text{ if }s\leq 2\\ (1,0,\ldots,0,2,4,\ldots,2^{n-x-1}),&\text{ else}.\end{cases}$
Proof.
Let $\mathcal{A}=(a_{0},\ldots,a_{n-1})$ be a basis of $S$ such that $M_{\mathcal{A}}(S)=M$. First note that $m_{0,0}=0$ if $s\leq 2$ (which means that $q(a_{0},a_{0})=0$) and $m_{0,0}=1$ else. By definition of the normal quaternionic matrix, $\mathcal{A}$ is chosen such that we get the maximum number of ’leading zeros’, i.e. $x^{\prime}\in\mathbb{N}_{0}$ maximal such that $q(a_{0},a_{i})=0$ for all
$i\in\begin{cases}\{0,\ldots,x^{\prime}-1\}&\text{if }s\leq 2\\ \{1,\ldots,x^{\prime}\}&\text{if }s>2.\end{cases}$
This means that $a_{0},\ldots,a_{x^{\prime}-1}$ resp. $a_{1},\ldots,a_{x^{\prime}}$ form a basis of $V(a_{0})=\{b\in G\mid q(a_{0},b)=0\}$. By definition, we get $x^{\prime}=x$. Define
$J=\begin{cases}\{x,x+1,\ldots,n-1\}&\text{if }s\leq 2\\ \{0,x+1,x+2,\ldots,n-1\}&\text{if }s>2.\end{cases}$
The values $(q(a_{0},a_{j})\mid j\in J)$ must be independent for else there would be a subset $I\subset J$ such that
[…]
and thus $\prod_{i\in I}a_{i}\in V(a_{0})$, a contradiction. Hence, the first row of $M$ is of the described form. ∎
As an immediate consequence of the above proof, we can specify exactly which bases of $S$ produce a minimal first row:
Corollary 5.2.
Let $M$ be a normal quaternionic matrix of order $n$ and type $\epsilon\in\{0,1\}$ with quaternionic structure $S:=S(M)$. Define $x$ as in 5.1. Let $\mathcal{A}=(a_{0},\ldots,a_{n-1})$ be any basis of $S$. The first row of $M_{\mathcal{A}}(S)$ is identical to that of $M$ if and only if
  • $s(S)\leq 2$ and $V(a_{0})=\operatorname{span}(a_{0},\ldots,a_{x-1})$ or
  • $s(S)>2$ and $V(a_{0})=\operatorname{span}(a_{1},\ldots,a_{x})$
Example 5.3.
Proposition 5.1 directly produces a list of ’potential’ normal quaternionic matrices of order $1$ and $2$ and the different matrices are clearly not isomorphic as the dimension of the value sets $V(a)$ is invariant under isomorphism. For type 00, we get exactly the matrices
$\begin{pmatrix}0\end{pmatrix},\quad\begin{pmatrix}0&0\\ 0&0\end{pmatrix},\quad\begin{pmatrix}0&1\\ 1&0\end{pmatrix}.$
which are the normal quaternionic matrices of $\mathbb{L}_{1,0}$, $\mathbb{L}_{1,0}^{2}$ and $\mathbb{L}_{1,0}\Delta$.
For type 1, we get the matrices
$\begin{pmatrix}0\end{pmatrix},\quad\begin{pmatrix}1\end{pmatrix},\quad\begin{pmatrix}0&0\\ 0&0\end{pmatrix},\quad\begin{pmatrix}0&1\\ 1&1\end{pmatrix},\quad\begin{pmatrix}1&0\\ 0&0\end{pmatrix},\quad\begin{pmatrix}1&2\\ 2&2\end{pmatrix}.$
One readily checks that these are the normal quaternionic matrices of (in the same order)
$\mathbb{L}_{1,1},\quad\mathbb{L}_{1},\quad\mathbb{L}_{1,1}^{2},\quad\mathbb{L}_{1,1}\Delta,\quad\mathbb{L}_{1,1}\times\mathbb{L}_{1},\quad\mathbb{L}_{1}^{2}(\cong\mathbb{L}_{1}\Delta).$
Comparing with the number of elementary type structures of order $1$ and $2$, this verifies the ETC for structures order $1$ and $2$ (also cf. Example 4.1).
To also characterize the structure of the second row, we introduce a few definitions and a somewhat technical lemma that will also be of use for applying our algorithm in the final section. Recall that we apply the lexicographic order on matrices row by row. We are often interested in all entries of a quaternionic matrix up to (with respect to that order) a certain row/column index.
Definition.
Given $x,y\in\mathbb{N}$ we say that $y$ is a factor of $x$ if $x*y<x$.
Let $M=(M_{i,j})_{0\leq i,j\leq n-1}$ be a quaternionic matrix of order $n\in\mathbb{N}$. Take $i,j,k,\ell\in\{0,\ldots,n-1\}$ with $(i,j)\leq_{\text{lex}}(k,\ell)$. We say that $M_{i,j}$ is a unique value (of $M$) up to $(k,\ell)$ if $M_{i,j}$ is a power of 2 that is no factor of any of the elements of $\{M_{i^{\prime},j^{\prime}}\mid i^{\prime}<j^{\prime},\ (i^{\prime},j^{\prime})\not=(i,j)\text{ and }(i^{\prime},j^{\prime})<_{\text{lex}}(k,\ell)\}$
Lemma 5.4.
Let $M$ be a normal quaternionic matrix of order $n$ and $i,j\in\{0,\ldots,n-1\}$ with $i<j$. The following hold:
  1. (a)
    If there is $k<i$ such that $M_{k,j}\not=0$ and for all $\ell<j$ with $\ell\not=i$, we have $M_{k,l}=0$ or $M_{i,\ell}$ is unique up to $(i,j)$, then $M_{k,j}$ is no factor of $M_{i,j}$.
  2. (b)
    If there is $\ell<j$ such that $M_{i,\ell}\not=0$ and for all $k<i$, we have $M_{k,\ell}=0$ or $M_{k,j}$ is unique up to $(i,j)$, then $M_{i,\ell}$ is no factor of $M_{i,j}$.
  3. (c)
    If $j+1>i$ and $M_{k,j}=0$ for all $k<i$, then $M_{i,j}\geq M_{i,j-1}$.
  4. (d)
    If there is $k<i$ such that for all $\ell\geq j$, we have $M_{k,\ell}\not\in\operatorname{span}(M_{k,0},\ldots,M_{k,\ell-1})$, then the following hold:
    1. (i)
      If $M_{i,j}\in\operatorname{span}(M_{k,j^{\prime}}\mid j^{\prime}<j)$, then $M_{i,j}\in\operatorname{span}(M_{i,j^{\prime}}\mid j^{\prime}<j)$
    2. (ii)
      If $M_{k,j}\in\operatorname{span}(M_{i,j^{\prime}}\mid j^{\prime}<j)$, then $M_{i,j}\in\operatorname{span}(M_{i,j^{\prime}}\mid j^{\prime}<j)$.
Proof.
(a): We apply the transformation $(i\to k\ast i)$ to obtain the quaternionic matrix $M^{\prime}$ of $S$ and set $M^{\prime\prime}:=M^{\prime}_{\text{red}}$. We clearly have $M^{\prime}_{i^{\prime},j^{\prime}}=M_{i^{\prime},j^{\prime}}$ if $i\not\in\{i^{\prime},j^{\prime}\}$. For $\ell\not=i$, $\ell<j$, we get
$M^{\prime}_{i,\ell}=M_{i,\ell}\ast M_{k,\ell}.$
If $M_{k,\ell}=0$, then $M^{\prime}_{i,\ell}=M_{i,\ell}$. Otherwise, $M_{i,\ell}$ is unique up to $(i,j)$ and thus, we still get
$M^{\prime}_{i,\ell}\not\in\operatorname{span}(M^{\prime}_{i^{\prime},j^{\prime}}\mid(i^{\prime},j^{\prime})<_{\text{lex}}(i,\ell)\}.$
Hence, we have $M^{\prime\prime}_{i,\ell}=M_{i,\ell}$ in either case.
Although we transformed the (implicitly employed) quaternionic basis by considering the reduced matrix $M^{\prime\prime}=M^{\prime}_{\text{red}}$, we still get $M^{\prime\prime}_{i^{\prime},j^{\prime}}=M_{i^{\prime},j^{\prime}}$ for all $(i^{\prime},j^{\prime})<_{\text{lex}}(i,j)$ since the exchanged basis quaternions are no factors of any other entries of $M$. As $M$ is normal, it follows that $M^{\prime\prime}_{i,j}\geq M_{i,j}$ and since $M^{\prime}_{i,j}=M_{i,j}\ast M_{k,j}$, we have that $M_{k,j}$ is no factor of $M_{i,j}$.
(b): This follows in the same way as (a), now considering the transformation $(j\to j\ast\ell)$.
(c): This follows similarly to (a) and (b), without the need to consider unique values, by the transformation $(j-1\leftrightarrow j)$.
(d): It follows immediately from the given condition that $\overline{M}_{2^{k},x}\not=\overline{M}_{2^{k},y}$ for $x<2^{j}\leq y$. By (b)(M4) and the definition of the completion, if
$M_{i,j}\in\operatorname{span}(M_{k,j^{\prime}}\mid j^{\prime}<j)=\{\overline{M}_{2^{k},x}\mid x\in\{0,\ldots,2^{j}-1\}\}$
then we already have $M_{i,j}\in\operatorname{span}(M_{i,j^{\prime}}\mid j^{\prime}<j)$. This shows (d)(i).
If now $M_{k,j}\in\operatorname{span}(M_{i,j^{\prime}}\mid j^{\prime}<j)$, we can apply (b)(M4) again, making use of the fact that $\overline{M}_{2^{k},x}\not=M_{k,j}$ for $x<2^{j}$ and also for $x\geq 2^{j+1}$ since $M_{k,j+1},\ldots,M_{k,n-1}$ are all independent from the previous entries of the row. Hence, there must be $y$ with $2^{j}\leq y<2^{j+1}$ such that $\overline{M}_{2^{i},y}=M_{k,j}$ which proves (d)(ii). ∎
Proposition 5.5.
Let $n\in\mathbb{N}$, $n\geq 3$ and $M\in\operatorname{Mat}_{n}(\mathbb{N}_{0})$ be a normal quaternionic matrix with level $s:=s(S(M))$. Then there are unique $d,r,t,u,v\in\mathbb{N}_{0}$ satisfying $t\leq u\leq d\leq v$ such that the first two rows of $M$ can be written as
$\left(\begin{array}[]{ccccccc}\begin{matrix}M_{0,0}&M_{0,1}\\ M_{1,0}&M_{1,1}\end{matrix}&B_{1}&B_{2}&B_{3}&B_{4}&B_{5}&B_{6}\end{array}\right)$
with matrices $B_{i}\in\operatorname{Mat}_{2\times k_{i}}(\mathbb{N}_{0})$ and $(k_{1},\ldots,k_{6})=(n-r-v-2,r,v-d,t,u-t,d-u)$, the case of zero columns being allowed and interpreted as a non existent block. If $s\in\{1,2\}$ and $M_{0,1}=0$, these matrices can be written as the following:
$B_{1}=\begin{pmatrix}0&\ldots&0\\ 0&\ldots&0\end{pmatrix},\quad B_{2}=\begin{pmatrix}0&\ldots&0\\ 2^{0}&\ldots&2^{r-1}\end{pmatrix},\quad B_{3}=\begin{pmatrix}0&\ldots&0\\ 2^{d}&\ldots&2^{v-1}\end{pmatrix}$
$B_{4}=\begin{pmatrix}2^{0}&\ldots&2^{t-1}\\ 0&\ldots&0\end{pmatrix},\quad B_{5}=\begin{pmatrix}2^{t}&\ldots&2^{u-1}\\ 2^{t}&\ldots&2^{u-1}\end{pmatrix},\quad B_{6}=\begin{pmatrix}2^{u}&\ldots&2^{d-1}\\ 2^{v}&\ldots&2^{v+d-u-1}\end{pmatrix}$
If $s>2$ and $M_{0,1}=0$, the entries of $B_{3},\ldots,B_{6}$ are multiplied by $2$ and the entries of $B_{2}$ are multiplied by a factor $f\in\{1,2\}$.
If $M_{0,1}\not=0$, the blocks $B_{1},B_{2}$ and $B_{3}$ do not appear while the entries of $B_{4},B_{5}$ and $B_{6}$ are multiplied by $2M_{0,1}$.
Proof.
By Proposition 5.1, the first row of $M$ has the asserted shape, setting $d:=n-x$ if $s\leq 2$ and $d:=n-x-1$ otherwise.
We write $S:=S(M)=(G,-1,q)$ and fix a basis $\mathcal{A}=(a_{0},\ldots,a_{n-1})$ of $S$ such that $M_{\mathcal{A}}(S)=M$. To keep notation brief, we usually only write $q$ for the quaternionic map instead of $q_{\mathcal{B}}$ for a basis $\mathcal{B}$ of the full 2B-group $B$ group whenever no confusion is imminent.
By (b)(M2) and (b)(M3), the entries $M_{0,1}$ and $M_{1,1}$ are of the desired shape. Let now $k\in\{2,\ldots,n-1\}$ be such that $M_{1,0},\ldots,M_{1,{k-1}}$ are as claimed. Set
$\alpha:=\operatorname{min}\{2^{x}\mid 2^{x}>M_{i,j}\text{ for all }(i,j)<_{\text{lex}}(1,k)\}.$
We have $M_{1,k}=\alpha$ if and only if $M_{1,k}$ is independent from the previous entries of $M$. Otherwise, $M_{1,k}$ is the quaternionic product of certain previously appearing entries. By Lemma 5.4(b) and induction hypothesis, the nonzero elements among $M_{1,2},\ldots,M_{1,k-1}$ are no factors of $M_{1,k}$. Furthermore, the entries $M_{1,0}$ and $M_{1,1}$ naturally appear in the first row. Hence, in this case, we can write
[…]
for some set $I\subset\{0,n-d,n-d+1,\ldots,n-1\}$.
In the following, we discern cases regarding in which of the blocks from $B_{1},\ldots,B_{6}$ the entry $M_{1,k-1}$ lies. We then show that in every case, $M_{1,k}$ either continues this block in the claimed way or begins a block with higher index. We first only consider the case $M_{0,1}=0$ although it does not make much difference for the arguments, as pointed out at the end of the proof.
Case 1: $M_{0,k}=0$
Case 1.1: $k=2$ or $M_{1,k-1}$ lies in $B_{1}$ (i.e. $M_{0,k-1}=M_{1,k-1}=0$)
[…]
$\left(a_{0},\ldots,a_{n-d-1},\prod_{i\in I}a_{i}\right)$
is independent and can be extended to a basis $\mathcal{A}^{\prime}$ of $S$. By 5.2, its quaternionic matrix $M^{\prime}$ has the same first row as $M$ and clearly $M^{\prime}_{1,2}=\ldots=M^{\prime}_{1,k-1}=0$. By the minimality of $M$, we now get
$0<M_{1,k}\leq M^{\prime}_{1,k}=q\left(a_{0},\prod_{i\in I}a_{i}\right)=M^{\prime}_{0,n-d}=M_{0,n-d}$
and thus $M_{1,k}=1$ if $s\leq 2$ and $M_{1,k}=2$ else. Note that the latter case begins $B_{2}$ multiplied by $f=2$ while the former case satisfies $f=1$ as asserted.
Case 1.2: $k>2$ and $M_{1,k-1}$ lies in $B_{2}$ or $B_{3}$
By Lemma 5.4(c), we have $M_{1,k}\geq M_{1,k-1}>0$ and as $M_{1,k-1}$ is a power of 2 and no factor of $M_{1,k}$, even $M_{1,k}\geq 2\cdot M_{1,k-1}$. If $M_{1,k}=\alpha$ we are done as in the previous case. In particular, this occurs if $M_{1,k-1}$ lies in $B_{3}$ since then $\alpha=2M_{1,k-1}$.
Otherwise, since $M_{1,k}\geq 2M_{1,k-1}$, there is $n-d\leq i\in I$ such that $M_{0,i}\not\in\{M_{1,k^{\prime}}\mid k^{\prime}<k\}$. Choose $\ell$ such that $M_{0,\ell}=2M_{1,k-1}$. Then $\ell\geq n-d$ is minimal such that $M_{0,\ell}\notin\{M_{1,k^{\prime}}\mid k^{\prime}<k\}$. This means that in particular $i\geq\ell$ and
$\left(a_{0},\ldots,a_{\ell-1},\prod_{i\in I}a_{i}\right)$
is independent and can be extended to a basis $\mathcal{A}^{\prime}$ of $S$. By 5.2, the matrix $M^{\prime}:=M_{\mathcal{A}^{\prime}}(S)$ has the same first row as $M$ and since $(a_{0},\ldots,a_{\ell-1})$ were not changed in the basis, we get $M_{1,k^{\prime}}=M^{\prime}_{1,k^{\prime}}$ for all $k^{\prime}<k$: This is clear for entries in $B_{1}$ or $B_{3}$, while for $B_{2}$ we point out that we have $M_{1,k^{\prime}}=M_{0,i}$ for some $i<\ell$ by the way we chose $\ell$. By the minimality of $M$, we get
$M_{1,k}\leq M^{\prime}_{1,k}=M^{\prime}_{0,\ell}=M_{0,\ell}=2M_{1,k-1},$
completing the case.
Case 2: $M_{0,k}\not=0$
[…]
$I=\emptyset\quad\text{or}\quad I\not\subset\{0,\ldots,k-1\}.$
Considering the transformation $(k-1\leftrightarrow k)$ in a similar way as in Lemma 5.4(c), the case $I=\emptyset$ (beginning or continuing $B_{4}$) can only occur if $M_{1,k-1}$ does not lie in $B_{5}$ or $B_{6}$.
It remains to consider the case $I\not\subset\{0,\ldots,k-1\}$, meaning that
$\left(a_{0},\ldots,a_{k-1},\prod_{i\in I}a_{i}\right)$
is independent and can be extended to a basis of $S$. As $M_{0,k},\ldots,M_{0,n-1}$ are unique up to $(1,k)$ by induction hypothesis, it now follows just like in Case 1.2 that $M_{1,k}=M_{0,k}$, completing Case 2.
Lastly, if $M_{0,1}\not=0$, by Proposition 5.1 we have
$M_{0,1}=\begin{cases}1,&\text{if $s\leq 2$}\\ 2,&\text{if $s>2$.}\end{cases}$
In either case, we get $M_{0,k}\not=0$ for all $k\geq 2$, so the blocks $B_{1},B_{2},B_{3}$ do not appear and we only have to consider case 2 above which works in the same way here. The only difference is that the first appearing block starts with entry $M_{0,2}=2M_{0,1}\in\{2,4\}$ as its first entry which carries forward as a factor for all other entries.
Note that the uniqueness of the parameters $d,f,r,t,u,v$ follows immediately from the uniqueness of the normal quaternionic matrix. ∎
We note the following specifications:
Corollary 5.6.
Let $M\in\operatorname{Mat}_{n}(\mathbb{N}_{0})$ be a normal quaternionic matrix and define $s,d,f,r,t,u,v$ as in Proposition 5.5. The following hold:
  1. (a)
    If $s\leq 2$ or $f=2$, we have $t\geq r$. Otherwise, we have $t\geq r-1$.
  2. (b)
    If $s\leq 2$, we have $2t\geq r+u$.
  3. (c)
    If $s=1$ and $d\not=n-1$, we have $t\leq r+v-d$.
Proof.
We keep the notation from the previous proof.
(a): If for $k\geq 2$, the entry $0\not=M_{0,k}$ appears as $M_{1,\ell}$ for $\ell<k$, we get $M_{1,k}\in\operatorname{span}(M_{1,k^{\prime}}\mid k^{\prime}<k)$ by Lemma 5.4(d)(d)(i). Applying Lemma 5.4(b), we can exclude $M_{1,\ell}=M_{0,k}$ as a factor. Hence, $M_{1,k}=0$ by Proposition 5.5.
As this happens for all second row entries of $B_{2}$ with the exception of $1$ if $s>2$ and $f=1$, the claim follows.
(b): Consider the transformation $(2\mapsto 1\ast 2)$: This does not change any of the entries at indices $(0,0)$, $(0,1)$, $(1,0)$, $(1,1)$ as $s\leq 2$ and clearly also not the entries in $B_{1}$, $B_{2}$ and $B_{3}$ while $B_{4}$ and $B_{5}$ are transformed into
$\begin{pmatrix}2^{0}&\ldots&2^{t-1}\\ 2^{0}&\ldots&2^{t-1}\end{pmatrix}\quad\text{and}\quad\begin{pmatrix}2^{t}&\ldots&2^{u-1}\\ 0&\ldots&0\end{pmatrix}.$
Applying Lemma 5.4(b) with the second row entries of $B_{2}$, we can further transform $B_{4}$ to
$\begin{pmatrix}2^{0}&\ldots&2^{r-1}&2^{r}&\ldots&2^{t-1}\\ 0&\ldots&0&2^{r}&\ldots&2^{t-1}\end{pmatrix}.$
Permuting the $u-t$ columns in front of the last $t-r$ columns of $B_{4}$, we can thus attain a block of the shape
$\begin{pmatrix}2^{0}&\ldots&2^{r+u-t}\\ 0&\ldots&0\end{pmatrix}$
instead of $B_{4}$ without changing the first row (cf. 5.2) or the first three blocks as seen above. By the minimality of $M$, we must have $t\geq r+u-t$.
(c): By Proposition 5.1, the dimension $\dim(V(a_{0}))=n-d$ is maximal. If $d\not=n-1$, it follows that $M_{0,0}=M_{1,0}=M_{0,1}=M_{1,1}=0$ and counting the number of zeroes in the second row with Proposition 5.5, we get $\dim(V(a_{1}))=n-r-v+t$. We thus have $n-r-v+t\leq n-d$ and the claim follows. ∎
Example 5.7.
Any normal quaternionic matrix of order up to 3 is uniquely defined by its first two rows and its type. There are exactly 17 isomorphy types of elementary type structures of order $3$ as can be seen in Lemma 2.9. The previous result gives rise to 4 options when the type is 0 and 14 options when the type is $1$. One of the latter does not define a quaternionic matrix as (b)(M4) is violated for
$M:=\begin{pmatrix}0&1&2\\ 1&1&0\\ 2&0&2\end{pmatrix}.$
This can be seen as the rows with index $3$ resp. $5$ of $\overline{M}$ are (0 1 0 1 2 3 2 3)01012323(0\ 1\ 0\ 1\ 2\ 3\ 2\ 3)( 0 1 0 1 2 3 2 3 ) and (0 2 1 3 0 2 1 3)02130213(0\ 2\ 1\ 3\ 0\ 2\ 1\ 3)( 0 2 1 3 0 2 1 3 ), so the entries $1$ and $2$ appear in both rows but do not appear in common columns.
The other 17 candidates from Proposition 5.5 thus have to be the normal quaternionic matrices of the 17 different isomorphy types of quaternionic structures of order 3 (cf. Lemma 2.9):
Type 0:
$\mathbb{L}_{1,0}^{3}:\begin{pmatrix}0&0&0\\ 0&0&0\\ 0&0&0\end{pmatrix},\ \mathbb{L}_{1,0}\times\mathbb{L}_{1,0}\Delta:\begin{pmatrix}0&0&0\\ 0&0&1\\ 0&1&0\end{pmatrix},\ \mathbb{L}_{1,0}^{2}\Delta:\begin{pmatrix}0&0&1\\ 0&0&2\\ 1&2&0\end{pmatrix},\ \mathbb{L}_{1,0}\Delta_{2}:\begin{pmatrix}0&1&2\\ 1&0&4\\ 2&4&0\end{pmatrix}.$
Type 1:
$\mathbb{L}_{1,1}^{3}:\begin{pmatrix}0&0&0\\ 0&0&0\\ 0&0&0\end{pmatrix},\ \mathbb{L}_{1,1}\times\mathbb{L}_{1,0}\Delta:\begin{pmatrix}0&0&0\\ 0&0&1\\ 0&1&0\end{pmatrix},\ \mathbb{L}_{1,1}\times\mathbb{L}_{1,1}\Delta:\begin{pmatrix}0&0&1\\ 0&0&0\\ 1&0&1\end{pmatrix},$
$\mathbb{L}_{1,1}^{2}\Delta:\begin{pmatrix}0&0&1\\ 0&0&2\\ 1&2&1\end{pmatrix},\ \mathbb{L}_{1,1}\Delta_{2}:\begin{pmatrix}0&1&2\\ 1&1&4\\ 2&4&2\end{pmatrix},\ \mathbb{L}_{3}:\begin{pmatrix}1&0&0\\ 0&0&1\\ 0&1&0\end{pmatrix},\ \mathbb{L}_{1,1}^{2}\times\mathbb{L}_{1}:\begin{pmatrix}1&0&0\\ 0&0&0\\ 0&0&0\end{pmatrix},$
$\mathbb{L}_{1}\times\mathbb{L}_{1,0}\Delta:\begin{pmatrix}1&0&0\\ 0&0&2\\ 0&2&0\end{pmatrix},\ \mathbb{L}_{1,1}\times\mathbb{L}_{1}^{2}:\begin{pmatrix}1&0&2\\ 0&0&0\\ 2&0&2\end{pmatrix},\ \mathbb{L}_{1}\times\mathbb{L}_{1,1}\Delta:\begin{pmatrix}1&0&2\\ 0&0&2\\ 2&2&2\end{pmatrix},$
$(\mathbb{L}_{1,1}\times\mathbb{L}_{1})\Delta:\begin{pmatrix}1&0&2\\ 0&0&4\\ 2&4&2\end{pmatrix},\ \mathbb{L}_{1}^{3}:\begin{pmatrix}1&2&4\\ 2&2&0\\ 4&0&4\end{pmatrix},\ \mathbb{L}_{1}^{2}\Delta:\begin{pmatrix}1&2&4\\ 2&2&8\\ 4&8&4\end{pmatrix}.$
Remark 5.8.
Given $n\in\mathbb{N}$, the number $e(n)$ of elementary type structures of order $n$ (cf. 2.9) gives a lower bound on the number of normal quaternionic structures. In order to show that the ETC holds for order $n$, it thus suffices to show that $e(n)$ also is an upper bound for the number of different normal quaternionic matrices of order $n$. Hence, we do not focus on giving a detailed and general description of normal quaternionic matrices of elementary type structures here. However, we point out a few details without proof. Furthermore, many examples for normal quaternionic matrices of order up to $6$ are included in Appendix A.
  1. (a)
    The matrices given in Example 4.1(b) are indeed the normal quaternionic matrices of the local type structures.
  2. (b)
    One can obtain the normal quaternionic matrix of a group extension $S\Delta$ from a normal quaternionic matrix of $S$ by adding an independent column and reducing. For example consider $S=\mathbb{L}_{1}\times(\mathbb{L}_{1}\times\mathbb{L}_{1})\Delta$ and $S\Delta$ with (cf. Example 4.7)
    $M(S)=\begin{pmatrix}1&2&4&8\\ 2&2&0&0\\ 4&0&4&16\\ 8&0&16&8\end{pmatrix}\quad\text{and}\quad M(S\Delta)=\begin{pmatrix}1&2&4&8&16\\ 2&2&0&0&32\\ 4&0&4&64&128\\ 8&0&64&8&256\\ 16&32&128&256&16\end{pmatrix}.$
  3. (c)
    For direct products, the situation is more involved. Take structures $S=(G,-1,q)$ and $S^{\prime}=(G^{\prime},-1,q^{\prime})$ with normal quaternionic matrices $M=M_{\mathcal{A}}(S)$ and $M^{\prime}=M_{\mathcal{A}^{\prime}}(S^{\prime})$ for bases $\mathcal{A}=(a_{0},\ldots,a_{n-1})$ of $S$ and $\mathcal{A}^{\prime}=(a_{0}^{\prime},\ldots,a_{n^{\prime}-1}^{\prime})$ of $S^{\prime}$. We conjecture that we get $M(S\times S^{\prime})=M_{\mathcal{C}}(S\times S^{\prime})$ for a basis $\mathcal{C}=(c_{0},\ldots,c_{n+n^{\prime}-1})$ that is close to a union of $\mathcal{A}$ and $\mathcal{A^{\prime}}$ in the following way: Consider $G$ and $G^{\prime}$ as subsets of $G\times G^{\prime}$ canonically. Then $c_{0}\in\{a_{0},a_{0}^{\prime},a_{0}a_{0}^{\prime}\}$ depending on the levels of $S$ and $S^{\prime}$ and
[…]
To close this section, we point out the following example to illustrate how the situation of direct products becomes slightly more involved as mentioned in Remark 5.8(c), especially when multiple factors have level different from one.
Example 5.9.
The situation for direct products also gives rise to quaternionic matrices of elementary type structures where nonzero entries appear that are no powers of 2. This is due to the fact that we have to fix $-1$ as the first basis element. For example, we get
$M(\mathbb{L}_{3})=\begin{pmatrix}1&0&0\\ 0&0&1\\ 0&1&0\end{pmatrix}\quad\text{and}\quad M(\mathbb{L}_{3}^{2})=\begin{pmatrix}1&0&0&0&0&2\\ 0&0&0&0&2&0\\ 0&0&0&3&0&0\\ 0&0&3&0&0&0\\ 0&2&0&0&0&0\\ 2&0&0&0&0&2\end{pmatrix}.$
To obtain this matrix, take the bases $(a_{0},a_{1},a_{2})$ and $(a^{\prime}_{0},a^{\prime}_{1},a^{\prime}_{2})$ of the two copies of $\mathbb{L}_{3}$ defining the normal quaternionic matrix of $\mathbb{L}_{3}$ and consider the basis $(a_{0}a_{0}^{\prime},a_{1},a^{\prime}_{1},a^{\prime}_{2},a_{2},a_{0})$ of $\mathbb{L}_{3}^{2}$.
For considering direct products where at least one factor has level $1$, the situation usually becomes simpler. For example consider the structure $\mathbb{L}_{1,0}\Delta\times\mathbb{L}_{3}$ with
$M_{\mathbb{L}_{1,0}\Delta}=\begin{pmatrix}0&1\\ 1&0\end{pmatrix}\quad\text{and}\quad M_{\mathbb{L}_{1,0}\Delta\times\mathbb{L}_{3}}=\begin{pmatrix}1&0&0&0&0\\ 0&0&0&0&1\\ 0&0&0&2&0\\ 0&0&2&0&0\\ 0&1&0&0&0\end{pmatrix}.$
6. The Computation
In this section, we present the algorithm applied to compute quaternionic matrices along with runtimes for some cases.
We start with the following reduction technique. Once the ETC has been confirmed for order up to $n-1$, in order to verify the ETC for order $n$, it is enough to show the ETC for nondegenerate structures that are neither rigid nor pythagorean by Proposition 2.8(a) and (d). Furthermore, depending on the level, we can also exclude some $2$-rigid cases. To make use of this, we include the following results that show how how to read these properties off normal quaternionic matrices.
Corollary 6.1.
Let $S$ be a quaternionic structure of order $n$ with normal quaternionic matrix $M$. Then $S$ is nondegenerate and neither rigid nor pythagorean if and only if every row of $\overline{M}$ has at least $4$ and at most $2^{n-1}$ zeroes.
In this case, we have $M_{0,1}=M_{1,0}=M_{1,1}=0$.
Proof.
The first part follows immediately from the definitions as the zeroes in a row of $\overline{M}$ correspond $1-1$ to the elements of $V(a)$ for some $a$. Considering Proposition 5.1 along with (b)(M2) and (b)(M3), we get the second part of the assertion. ∎
Lemma 6.2.
Let $S=(G,-1,q)$ be a nondegenerate quaternionic structure of order $n$ with normal quaternionic matrix $M$. We have $s(S)=1$ if and only if $M$ has only zeroes on the diagonal. Else, we have $s(S)=2$ if and only if $M_{0,0}=0$. Lastly, we get $s(S)=4$ if and only if $M_{0,0}=1=\overline{M}_{I,J}$ for some $I,J\subset\{1,\ldots,n-d-2\}$ where $d$ is defined in the same way as in Proposition 5.5, i.e. $\overline{M}_{\{0\},I}=\overline{M}_{\{0\},J}=0$.
Proof.
The statements about level $\leq 2$ are an immediate consequence of (b)(M3) and the fact that
$-1\in D(\langle 1,1\rangle)\iff q(-1,-1)=0\iff M_{0,0}=0.$
If $s(S)=4$, we have $M_{0,0}=1$ as seen before and $-1\in D(\langle 1,1,1,1\rangle)$. Note that by the correspondence of $q$ and $M$ (cf. 4.4) the existence of $I,J$ as asserted above is equivalent to the existence of $a,b\in G$ such that $q(a,b)=1$ and $q(-1,a)=q(-1,b)=0$. By a classical property which also holds in the abstract setting (cf. [Marshall_AbstractWittRings, Proposition 2.10]), we have
$D(\langle 1,1,1,1\rangle)=\bigcup_{a\in D(\langle 1,1,1\rangle}D(\langle 1,a\rangle)=\bigcup_{a\in D(\langle 1,1,1\rangle)}\{x\mid q(-a,x)=0\},$
$D(\langle 1,1,1\rangle)=\bigcup_{b\in D(\langle 1,1\rangle)}D(\langle 1,b\rangle)=\bigcup_{\{b\mid q(-1,b)=0\}}\{y\mid q(-b,y)=0\}.$
Hence, we find $a\in D(\langle 1,1,1\rangle)$ and $b\in D(\langle 1,1\rangle)$ such that $q(-a,-1)=q(-b,a)=q(-1,b)=0$. By bimultiplicativity and symmetry of $q$, and applying $q(-1,-1)=1$, it follows that $q(-1,a)=1$ and $q(a,b)=q(a,-1)\ast q(a,-b)=1$. As pointed out above, the existence of $a,b$ proves the existence of $I,J$ as asserted. Conversely, if there are $a,b\in G$ such that $q(a,b)=1$ and $q(-1,a)=q(-1,b)=0$, we can do the same computations as above to show that $-1\in D(\langle 1,a\rangle)$ and $a\in D(\langle 1,1,1\rangle)$, completing the proof. ∎
Remark 6.3.
The previous two results enable us to apply the results stated in Proposition 2.8 to exclude some cases. We explicitly collect these here for a better overview. To confirm the ETC for order $n$, whenever we know that the ETC holds for order up to $n-1$, we can exclude quaternionic structures whose normal quaternionic matrix $M$ has one of the following properties:
  • There is a row of the completion $\overline{M}$ that has only zeroes or fewer than $4$ zeroes.
  • We have $M_{n-1,n-1}=0$ (i.e. level $1$) and any row of the completion has fewer than $8$ zeroes.
  • We have $M_{0,0}$ (i.e. level at most $2$) and any row of the completion starting with 00 has fewer than $8$ zeroes.
  • We have $M_{0,0}=M_{1,2}=1$ and $M_{0,3}=2$ or $n=3$. This means that $-1$ is $2$-rigid and we have level $4$, corresponding to the parameters (cf. Proposition 5.5) $d=n-3$, $f=1$, $r=1$ arbitrary $v,t,u$. Note that in this case, we always have $v=n-2$ by the restrictions presented in Corollary 5.6.
Our computations show the following. For computation times, we refer to Remark 6.6.
Theorem 6.4.
  1. (a)
    All quaternionic structures of order up to $6$ are of elementary type.
  2. (b)
    All quaternionic structures of order $7$ and level in $\{1,2\}$ are of elementary type. Furthermore all those of order $7$ and level at least $4$ with parameters (cf. Proposition 5.5)
    $(d,f,r,v,t,u)\not\in\Big{\{}(3,1,0,3,0,0),\quad(4,1,0,4,1,1),\quad(4,1,0,4,2,2),\quad(4,2,1,4,1,1)\Big{\}}$
    are of elementary type.
Proof.
(a): By the results of Carson and Marshall as cited in 2.8(c), every quaternionic structure of order up to $5$ is of elementary type. Hence, it suffices to compute the correct upper bound for the nonpythagorean, nonrigid and nondegenerate normal quaternionic matrices, also excluding the existence of $2$-rigid elements in the cases mentioned in Lemma 2.9. By Remark 2.10 there are $81=\sum_{k>1}e_{k}(6)$ nonrigid and nondegenerate normal quaternionic matrices of order $6$. One readily checks that among those, there are $10$ of level $1$, $28$ of level $2$ and $43$ of level at least $4$.
By Proposition 2.8, we can further exclude all $2$-rigid structures of level $1$, all structures of level $2$ where an element in $D(\langle 1,1\rangle)$ is $2$-rigid and all structures of level $4$ where $1$ is a $2$-rigid element. Regarding level $1$, there are $7$ strictly $2$-rigid nondegenerate structures since these are exactly the direct products of two group extensions, cf. proof of Lemma 2.9. With similar counting methods, one obtains that among the 28 structures with level $2$, nine have a $2$-rigid element in $D(\langle 1,1\rangle)$. Regarding level $4$, there are no further exclusions since the elementary type structures of level $4$ where $1$ is $2$-rigid are exactly $\mathbb{L}_{3}$ of order $3$ and its iterated group extensions which are rigid.
Applying this, we verified the ETC for order $6$ by running Algorithm 1 with the above restrictions, computing that there are at most $4$ resp. $19$ resp. $43$ normal quaternionic matrices of order $6$ and level $1$ resp. $2$ resp. at least $4$.
The argument for (b) is the same, building further on the results of part (a). Here, the lower bound on the number of normal quaternionic matrices given by the elementary type structures is $4$ resp. $65$ resp. $199$ for level $1$ resp. $2$ resp. at least $4$. For level $1$ and $2$, these numbers are obtained as upper bounds by Algorithm 1. For level $4$, excluding the parameters
$(3,1,0,3,0,0),\quad(4,1,0,4,1,1),\quad(4,1,0,4,2,2),\quad(4,2,1,4,1,1),$
we obtain $198$ normal quaternionic matrices which verifies the ETC in these cases since there is one elementary type structure with parameters $(3,1,0,3,0,0)$, namely $\mathbb{L}_{1,0}\Delta\times(\mathbb{L}_{1,1}\times\mathbb{L}_{1})\Delta_{3}$, and no elementary type structures for the other excluded parameters. ∎
We briefly outline some pseudo code in Algorithm 1 to classify normal quaternion matrices while the complete version we used can be found via the following link: https://github.com/nicolo314/NormalQuaternionicMatrices.
In the program, partial quaternionic matrices (or what is aimed to be completed to such) are created row by row and via depth search in order to minimize the data that has to be stored simultaneously. To keep track of the current object, we save our current data in an $n\times n$ matrix $M$ whose entries are tuples. The first entry of the tuple in $M(i,j)$ is a partial quaternionic matrix $A$ with entries up to position $(i,j)$ and the second entry of $M(i,j)$ is a list of possible extensions at index $(i,j+1)$ of $A$ that still have to be handled. We will make use of functions that increase resp. decrease $(i,j)$ to the next resp. last index, taking into considerations that elements of a quaternionic matrix up to the diagonal of a given row are already determined by previous entries.
Quaternionic matrices are saved as a class object, among their attributes are a partial completion, the unique values up to the last computed element and their parameters. Whenever a transformation is found that produces a smaller reduced matrix, or a violation of the common slot property (b)(M4) or another assumption (in particular: rigidity) is detected, the current class object is discarded. This will be called validity check in the following pseudo code, a function returning either true or false depending on whether the object under consideration is valid or not respectively. Whenever an object is discarded, the program continues with the last object that has not yet been discarded. Note that we only consider selected transformations as the exponentially growing number of possible transformations takes up too much computation time to be practical. As explained before in the proof of Theorem 6.4, it suffices that the resulting list contains all normal quaternionic matrices of the considered properties and the number of those matrices is exactly the number of elementary types with those properties. We discern the cases of quaternionic matrices with level $1$, level $2$, and level at least $4$. To properly deal with this distinction, for $s_{\text{range}}\subseteq\{1,2,4\}$, we define
$\overline{s_{\text{range}}}:=\begin{cases}s_{\text{range}}&\text{ if }4\not\in s_{\text{range}}\\ s_{\text{range}}\cup\{2^{k}\mid k\geq 2\}\cup\{\infty\}&\text{ else.}\end{cases}$
Algorithm 1 Classification of Quaternionic Structures
$n\in\mathbb{N}_{\geq 2}$, $s_{\text{range}}\subseteq\{1,2,4\}$
List $L$ containing all nondegenerate, nonpythagorean and nonrigid normal quaternionic matrices of order $n$ and level in $\overline{s_{\text{range}}}$.
function Classify Quaternionic Structures($n,s_{\text{range}}$)
     $L=\emptyset$
     for all $s\in s_{\text{range}}$ do
         for all suitable parameters $p=(d,f,r,v,t,u)$ do
              $A\leftarrow$ partial quaternionic matrix in $M_{2,n}(\mathbb{N}_{0})$ with parameters $p$ (cf. Proposition 5.5)
              if Validity Check($A$) = true  then
                  $A\leftarrow$Add new row($A$)
                  $i\leftarrow 2,\quad j\leftarrow 2$
                  $M(i,j)\leftarrow(A,\,$possible next elements($A$)))))
                  while $(i,j)>_{\text{lex}}(1,n)$ do
                       $A\leftarrow M(i,j)[0],\quad X\leftarrow M(i,j)[1]$
                       if $X\neq\emptyset$ then
                           $A\leftarrow A$ appended by $x\in X$, delete $x$ from $X$
[…]
                                $M(i,j+1)\leftarrow A$
                                increase($(i,j)$)
                       else if $j=n-1$ then
                           if $i=n-1$ then
[…]
                                    $L\leftarrow L\cup\{A\}$
                                    Decrease($(i,j)$)
                           else
                                increase($(i,j)$)
                                $A\leftarrow$Add new row($A$)
                                $M(i,j)\leftarrow(A,\,$possible next elements($A$)))))
                       else
                           decrease($(i,j)$)
     $L\leftarrow\{A\in L\mid A\text{ is reduced}\}$
     return $L$
function Add new row($A$)
     if $A\in M_{k,n}(\mathbb{N}_{0})$ for some $k<n$ then
         return $A$ appended by partial row $k+1$ up to diagonal using (b)(M2) and (b)(M3)
function possible next elements($A$)
     if Validity Check($A$) = true  then
         if $A\in M_{k,n}(\mathbb{N}_{0})$ for some $k\leq n$ then
              return $\emptyset$
         else
              return set of possible next row elements considering Lemma 5.4
     else
         return $\emptyset$
Remark 6.5.
In this remark, we highlight some computational aspects.
  • As mentioned before, it was not possible to check all applicable transformations during the validity check to assure the minimality of the constructed quaternion matrix a priori. But since we have a lower bound on the number of abstract Witt rings of a given order by Remark 2.10, we could conclude the minimality of the quaternion matrices a posteriori by showing the effectiveness of the lower bound, implying that the computed matrices are in fact normal quaternion matrices.
  • While developing the code, we constantly extended the set of considered transformations by inspecting provisional results of earlier versions and identifying matrices in the output that are not normal quaternionic matrices. We further considered partial quaternionic matrices appearing as intermediate results to find additional transformations. By doing so, we established a list of transformations mainly consisting of products of few simple transformations.
  • Instead of applying transformations explicitly, we aimed to use transformations implicitly in order to exclude possible extensions of a given partial matrix as soon as possible, leading to a better performance. The theoretical background is given in Lemma 5.4, mainly in parts (a) and (b). As another example of implicit transformations, we implemented first row checks and second row checks. Since the parameters introduced in Proposition 5.5 are unique, we can check whether permuting the basis element corresponding to the current row to the beginning of the basis would result in another set of parameters yielding a lexicographically smaller matrix.
  • The common slot property (b)(M4) can in general be checked only for completed rows. Nevertheless, we implemented some rules to exclude possible next entries or to discard matrices early. These work along the lines of Lemma 5.4(d).
  • The (partial) completion was not saved as a matrix but as a list of dictionaries, in order to check the common slot property (b)(M4) efficiently. We save a dictionary for each row, and the entry for the key $k$ in dictionary $i$ contains a list of those $j$, for which $\overline{M}_{i,j}=k$. The common slot check for rows $i,i^{\prime}$ then boils down to showing that for each common key $k$ in the dictionaries belonging to $i,i^{\prime}$, the corresponding lists have nonempty intersection.
In the following remark we discuss the runtime and our approaches to shorten the runtime, which are based on parallelization techniques.
Remark 6.6.
  1. (a)
    In order to tap the full potential of a modern computer architecture with multiple cores, we parallelized some of our calculations. Mostly, we simply parallelized the different choices for parameters $s,d,f,r,t,u,v$, i.e. the different choices for the first two rows, as presented in Proposition 5.5. This works well if there are no parameters with tremendously large runtime, so in particular for $n\leq 6$.
  2. (b)
    For $n=7$ and certain sets of parameters, we used a more involved strategy that sped up the run time and enabled us to keep better track of intermediate results, but required multiple executions of the program with different inputs. This happened in particular for $n=7,s=2$ and the following three sets of parameters
    $p=(d,r,f,r,v,t,u)\in\Big{\{}(4,1,0,4,1,1),\quad(4,1,0,4,2,2),\quad(4,1,1,4,1,1)\Big{\}},$
    while all others parameters for $n=7,s=2$ could be computed within a few minutes each. We will explain our strategy for the case $p=(4,1,0,4,1,1)$ in detail:
    1. (1)
      First of all, we computed all extensions to $3\times 7$ matrices that could not be excluded by validity checks.
    2. (2)
      We then resumed our calculations for these $3\times 7$ matrices. For all such extensions except for
      $\begin{pmatrix}0&0&0&1&2&4&8\\ 0&0&0&0&16&32&64\\ 0&0&0&128&256&512&1024\end{pmatrix},$
      the program terminated relatively quickly. We handled these ’easy’ extensions in parallel.
    3. (3)
      For the $3\times 7$ matrix displayed above we computed the possible extensions up to the entry $(3,4)$. This means we added a partial fourth row, whose first 4 entries are already clear by the symmetry of a quaternionic matrix and computed one further element. We obtained the 10 possibilities
      $A_{0}=\begin{pmatrix}0&0&0&1&2&4&8\\ 0&0&0&0&16&32&64\\ 0&0&0&128&256&512&1024\\ 1&0&128&1&x\end{pmatrix}$
      where we had
[…]
      These cases were then again handled in parallel, where the cases for
[…]
  3. (c)
    We also applied this strategy for parameters for $n=7,s=4$, but used even more check points in between, at which we parallelized certain calculations. While it sometimes happens that matrices with high powers of 2 occurring early in the matrix are discarded quickly due to violation of rigidity constraints, most of the time, a high power of 2 occurring in the first two rows of the corresponding quaternionic matrix indicates a huge computational effort to be made.
[…]
    parameter runtime
[…]
    Table 3. Computed parameters with longest runtimes for $n=7$ and $s\geq 4$
    We were not yet able to produce results for the remaining $4$ choices of parameters
    $p\in\{(3,1,0,3,0,0),\quad(4,1,0,4,2,2),\quad(4,1,0,4,1,1),\quad(4,2,1,4,1,1)\}.$
[…]
  4. (d)
    Finally to have an overview consisting of comparable results, we include the runtimes for various cases in Table 4, where those for $n=7,s\in\{2,4\}$ are not obtained from a single run but as the sum of all partial executions, further assuming that no computations have been executed in parallel.
    Our actual implementation of Algorithm 1 was written in Python 3.12.4 and run on a MacBook Pro with Apple M4 Pro Chip and 48 GB RAM.
    $s$ $n$ 4 5 6 7
[…]
    Table 4. Runtimes
    Among the parameters for $n=7,s=1$, there was one parameter which was handled in about 7 minutes, one in about 3 minutes and all the others in less than a minute. For most of the parameters, the computations were done in only a few seconds.
  5. (e)
    While executing the computations for quaternionic structures of order $n=7$, we reached the limits of what is computable in reasonable time. We identified the following problems:
    • In most of the cases, high powers of 2 led to big additional effort. This comes from the obvious fact that given a partial quaternionic matrix in which $2^{0},\ldots 2^{k}$ occur, the next entry has to be chosen in $\{0,\ldots,2^{k+1}\}$. Hence introducing the next power of 2 results in doubling the possibilities for all upcoming entries of the matrix. It would be thus desirable to have stronger conditions on possible next entries, especially with respect to excluding 2-powers in the sense of Lemma 5.4.
    • So far, we do not have criteria to identify nonreduced matrices while computing them. Instead, we excluded them when already having complete matrices, so after the computational effort is already done. E.g. for $n=7,s=4$, we computed the following matrices, both of them determining the same quaternionic structure $\mathbb{L}_{1,0}\Delta\times(\mathbb{L}_{1}\times\mathbb{L}_{1,1}\Delta)\Delta_{2}$, but only the second one is reduced (also cf. Example 4.7):
      $\begin{pmatrix}1&0&0&0&2&4&8\\ 0&0&0&0&2&16&32\\ 0&0&0&64&0&0&0\\ 0&0&64&0&0&0&0\\ 2&2&0&0&2&33&128\\ 4&16&0&0&33&4&256\\ 8&32&0&0&128&256&8\end{pmatrix},\quad\begin{pmatrix}1&0&0&0&2&4&8\\ 0&0&0&0&2&16&32\\ 0&0&0&64&0&0&0\\ 0&0&64&0&0&0&0\\ 2&2&0&0&2&128&256\\ 4&16&0&0&128&4&512\\ 8&32&0&0&256&512&8\end{pmatrix}.$
      Even though the computation time for the nonreduced matrix is less than the time for the reduced matrix (mostly because it consists of smaller entries), it costs a lot of effort which would be desirable to exclude beforehand using a theoretical argument. These ’copies’ of isomorphic structures also seem to appear in intermediate results in a great number, even if they cannot be extended to a normal quaternionic matrix. This appears to be a major reason for the extreme computation times in certain cases.
    • From a purely computational point of view, it does not seem to be obvious a priori which parameters (or subcases thereof) need the most runtime. It appears that the most time consuming parameters usually have large $d$ and small values for $v-d$ but some parameters that satisfy these conditions could be handled with surprising ease. Additionally, while the ’critical’ parameters usually satisfy $r=0$, there was one exception for order $7$ and level $4$, namely $(4,2,1,4,1,1)$, which proved to be extremely difficult to handle. Criteria to identify these critical parameters and their most time consuming subcases would enable us to apply automated strategies for parallelization of calculations, instead of relying on manual inspections for each case.
We conclude this article with a comparison to the most recent classification result.
Remark 6.7.
As mentioned before, Carson and Marshall first computationally verified the ETC for order $5$ in their 1982 article [CarsonMarshall_Decomposition]. To compute all abstract Witt rings for a given group $G$ of exponent $2$, they consider the value groups $V(a)$ for $a\in G$. There are some similarities to the approach presented here, as these value groups can be recovered by considering the zeroes in the rows of the completion $\overline{M}$ of a given quaternionic matrix in the following way: if $a\in G$ is associated with row and column $k_{a}$ in the completion, the set $V(a)$ consists of all $b\in G$ such that $\overline{M}_{k_{a},k_{b}}=0$ for $b$ associated with row and column $k_{b}$.
While we were not able to recreate the computations done by Carson and Marshall due to sparse documentation, it appears that our possibilities for early validity checks as described before enhance efficiency. However, we cannot present any quantitative comparisons with the computation times achieved by Carson and Marshall.
Appendix A Lists of Normal Quaternionic Matrices
We include the list of all normal quaternionic matrices of all quaternionic structures of order $4$ and all those of orders $5$ and $6$ that are nondegenerate, nonpythagorean and nonrigid. Instead of displaying the entire matrices, we only include the upper right triangle (row by row, including the first diagonal entry) as this uniquely describes the quaternionic matrix (along with the level). For the same level, the matrices are sorted lexicographically.
normal quaternionic matrices of order 4
level 1:
$\mathbb{L}_{1,0}^{4}$ (0,0,0,0,0,0,0)
$\mathbb{L}_{1,0}^{2}\times\mathbb{L}_{1,0}\Delta$ (0,0,0,0,0,0,1)
$\mathbb{L}_{1,0}\times\mathbb{L}_{1,0}^{2}\Delta$ (0,0,0,0,0,1,2)
$\mathbb{L}_{1,0}\times\mathbb{L}_{1,0}\Delta_{2}$ (0,0,0,0,1,2,4)
$\mathbb{L}_{1,0}^{3}\Delta$ (0,0,0,1,0,2,4)
$\mathbb{L}_{4,0}$ (0,0,0,1,1,0,0)
$\mathbb{L}_{1,0}\Delta^{2}$ (0,0,0,1,2,0,0)
$(\mathbb{L}_{1,0}\times\mathbb{L}_{1,0}\Delta)\Delta$ (0,0,0,1,2,4,8)
$\mathbb{L}_{1,0}^{2}\Delta_{2}$ (0,0,1,2,4,8,16)
$\mathbb{L}_{1,0}\Delta_{3}$ (0,1,2,4,8,16,32)
level 2:
$\mathbb{L}_{1,1}^{4}$ (0,0,0,0,0,0,0)
$\mathbb{L}_{1,1}^{2}\times\mathbb{L}_{1,0}\Delta$ (0,0,0,0,0,0,1)
$\mathbb{L}_{1,1}\times\mathbb{L}_{1,0}^{2}\Delta$ (0,0,0,0,0,1,2)
$\mathbb{L}_{1,1}\times\mathbb{L}_{1,0}\Delta_{2}$ (0,0,0,0,1,2,4)
$\mathbb{L}_{1,1}^{2}\times\mathbb{L}_{1,1}\Delta$ (0,0,0,1,0,0,0)
$\mathbb{L}_{1,1}\times\mathbb{L}_{1,1}^{2}\Delta$ (0,0,0,1,0,0,2)
$\mathbb{L}_{1,1}^{3}\Delta$ (0,0,0,1,0,2,4)
$\mathbb{L}_{4,1}$ (0,0,0,1,1,0,0)
$\mathbb{L}_{1,0}\Delta\times\mathbb{L}_{1,1}\Delta$ (0,0,0,1,2,0,0)
$(\mathbb{L}_{1,1}\times\mathbb{L}_{1,0}\Delta)\Delta$ (0,0,0,1,2,4,8)
$\mathbb{L}_{1,1}\times\mathbb{L}_{1,1}\Delta_{2}$ (0,0,1,2,0,0,4)
$\mathbb{L}_{1,1}\Delta^{2}$ (0,0,1,2,0,2,0)
$(\mathbb{L}_{1,1}\!\!\times\!\!\mathbb{L}_{1,1}\Delta)\Delta$ (0,0,1,2,0,4,8)
$\mathbb{L}_{1,1}^{2}\Delta_{2}$ (0,0,1,2,4,8,16)
$\mathbb{L}_{1,1}\Delta_{3}$ (0,1,2,4,8,16,32)
level $>2$:
$\mathbb{L}_{1,1}^{3}\times\mathbb{L}_{1}$ (1,0,0,0,0,0,0)
$\mathbb{L}_{1,1}^{2}\times\mathbb{L}_{1}^{2}$ (1,0,0,2,0,0,0)
$\mathbb{L}_{1,1}\times\mathbb{L}_{1}^{3}$ (1,0,2,4,0,0,0)
$\mathbb{L}_{1,1}\times\mathbb{L}_{1}\times\mathbb{L}_{1,0}\Delta$ (1,0,0,0,0,0,2)
$\mathbb{L}_{1,1}\times\mathbb{L}_{1}\times\mathbb{L}_{1,1}\Delta$ (1,0,0,2,0,0,2)
$\mathbb{L}_{1,1}\times(\mathbb{L}_{1,1}\times\mathbb{L}_{1})\Delta$ (1,0,0,2,0,0,4)
$\mathbb{L}_{1,1}\times\mathbb{L}_{1}^{2}\Delta$ (1,0,2,4,0,0,8)
$\mathbb{L}_{1,1}\times\mathbb{L}_{3}$ (1,0,0,0,0,0,1)
$\mathbb{L}_{1}^{4}$ (1,2,4,8,0,0,0)
$\mathbb{L}_{1}^{2}\times\mathbb{L}_{1,0}\Delta$ (1,0,0,2,4,0,0)
$\mathbb{L}_{1}^{2}\times\mathbb{L}_{1,1}\Delta$ (1,0,2,4,0,4,0)
$\mathbb{L}_{1}\times\mathbb{L}_{1,0}^{2}\Delta$ (1,0,0,0,0,2,4)
$\mathbb{L}_{1}\times\mathbb{L}_{1,1}^{2}\Delta$ (1,0,0,2,0,2,4)
$\mathbb{L}_{1}\times(\mathbb{L}_{1,1}\times\mathbb{L}_{1})\Delta$ (1,0,2,4,0,4,8)
$\mathbb{L}_{1}\times\mathbb{L}_{1}^{2}\Delta$ (1,2,4,8,0,0,16)
$\mathbb{L}_{1}\times\mathbb{L}_{1,0}\Delta_{2}$ (1,0,0,0,2,4,8)
$\mathbb{L}_{1}\times\mathbb{L}_{1,1}\Delta_{2}$ (1,0,2,4,2,4,8)
$\mathbb{L}_{1}\times\mathbb{L}_{3}$ (1,0,0,2,2,0,0)
$(\mathbb{L}_{1,1}^{2}\times\mathbb{L}_{1})\Delta$ (1,0,0,2,0,4,8)
$(\mathbb{L}_{1,1}\times\mathbb{L}_{1}^{2}$ (1,0,2,4,0,8,16)
$\mathbb{L}_{1}^{3}\Delta$ (1,2,4,8,0,16,32)
$(\mathbb{L}_{1}\times\mathbb{L}_{1,0}\Delta)\Delta$ (1,0,0,2,4,8,16)
$(\mathbb{L}_{1}\times\mathbb{L}_{1,1}\Delta)\Delta$ (1,0,2,4,2,8,16)
$(\mathbb{L}_{1,1}\times\mathbb{L}_{1})\Delta_{2}$ (1,0,2,4,8,16,32)
$\mathbb{L}_{1}^{2}\Delta_{2}$ (1,2,4,8,16,32,64)
$\mathbb{L}_{3}\Delta$ (1,0,0,2,1,4,8)
normal quaternionic matrices of order 5 (nondegenerate, nonpythagorean, nonrigid) level 1: $\mathbb{L}_{1,0}\Delta\times\mathbb{L}_{1,0}^{2}\Delta$ (0,0,0,0,1,0,0,2,4,0,0) $\mathbb{L}_{1,0}\Delta\times\mathbb{L}_{1,0}\Delta_{2}$ (0,0,0,0,1,2,4,0,8,0,0) level 2: $\mathbb{L}_{1,0}\Delta\times\mathbb{L}_{1,1}^{2}\Delta$ (0,0,0,0,1,0,0,2,4,0,0) $\mathbb{L}_{1,1}\Delta\times\mathbb{L}_{1,0}^{2}\Delta$ (0,0,0,0,1,0,2,0,4,0,0) $\mathbb{L}_{1,1}\Delta\times\mathbb{L}_{1,0}\Delta_{2}$ (0,0,0,0,1,2,4,0,8,0,0) $\mathbb{L}_{1,1}\Delta\times\mathbb{L}_{1,1}^{2}\Delta$ (0,0,0,1,2,0,0,2,0,4,0) $\mathbb{L}_{1,0}\Delta\times\mathbb{L}_{1,1}\Delta_{2}$ (0,0,0,1,2,4,0,0,0,0,8) $\mathbb{L}_{1,1}\Delta\times\mathbb{L}_{1,1}\Delta_{2}$ (0,0,1,2,4,0,0,4,8,0,0) level $>2$: $\mathbb{L}_{5}$ (1,0,0,0,0,0,0,1,1,0,0) $\mathbb{L}_{1,0}\Delta\times\mathbb{L}_{3}$ (1,0,0,0,0,0,0,1,2,0,0) $\mathbb{L}_{1}\times\mathbb{L}_{4,0}$ (1,0,0,0,0,0,0,2,2,0,0) $\mathbb{L}_{1}\times\mathbb{L}_{1,0}\Delta^{2}$ (1,0,0,0,0,0,0,2,4,0,0) $\mathbb{L}_{1,1}\Delta\times\mathbb{L}_{3}$ (1,0,0,0,2,0,0,2,1,0,0) $\mathbb{L}_{1}\times\mathbb{L}_{4,1}$ (1,0,0,0,2,0,0,2,2,0,0) $\mathbb{L}_{1}\times\mathbb{L}_{1,0}\Delta\times\mathbb{L}_{1,1}\Delta$ (1,0,0,0,2,0,0,2,4,0,0) $\mathbb{L}_{1,0}\Delta\times(\mathbb{L}_{1,1}\times\mathbb{L}_{1})\Delta$ (1,0,0,0,2,0,0,4,8,0,0) $\mathbb{L}_{1}\times\mathbb{L}_{1,1}\Delta^{2}$ (1,0,0,2,4,0,0,4,2,0,0) $\mathbb{L}_{1,1}\Delta\times(\mathbb{L}_{1,1}\times\mathbb{L}_{1})\Delta$ (1,0,0,2,4,0,0,4,8,0,0) $\mathbb{L}_{1}^{2}\times\mathbb{L}_{3}$ (1,0,0,2,4,2,0,0,0,0,0)
normal quaternionic matrices of order 6, level 1
(nondegenerate, nonpythagorean, nonrigid) $\mathbb{L}_{1,0}\Delta\times\mathbb{L}_{1,0}^{3}\Delta$ (0,0,0,0,0,1,0,0,0,2,0,0,4,8,0,0) $\mathbb{L}_{1,0}^{2}\Delta^{2}$ (0,0,0,0,0,1,0,0,0,2,0,4,0,8,0,0) $\mathbb{L}_{1,0}^{2}\Delta\times\mathbb{L}_{1,0}\Delta_{2}$ (0,0,0,0,0,1,0,0,0,2,4,8,0,16,0,0) $\mathbb{L}_{6,0}$ (0,0,0,0,0,1,0,0,1,0,1,0,0,0,0,0) $\mathbb{L}_{1,0}\Delta\times\mathbb{L}_{4,0}$ (0,0,0,0,0,1,0,0,1,0,2,0,0,0,0,0) $\mathbb{L}_{1,0}\Delta^{3}$ (0,0,0,0,0,1,0,0,2,0,4,0,0,0,0,0) $\mathbb{L}_{1,0}\Delta\times(\mathbb{L}_{1,0}\times\mathbb{L}_{1,0}\Delta)\Delta$ (0,0,0,0,0,1,0,0,2,0,4,0,8,0,16,0) $\mathbb{L}_{1,0}\Delta\times\mathbb{L}_{1,0}^{2}\Delta_{2}$ (0,0,0,0,0,1,0,2,4,0,8,16,0,32,0,0) $\mathbb{L}_{1,0}\Delta\times\mathbb{L}_{1,0}\Delta_{3}$ (0,0,0,0,0,1,2,4,8,0,16,32,0,64,0,0) $\mathbb{L}_{1,0}\Delta_{2}^{2}$ (0,0,0,0,1,2,4,8,0,0,16,0,0,0,0,32)
normal quaternionic matrices of order 6, level 2
(nondegenerate, nonpythagorean, nonrigid) $\mathbb{L}_{1,0}\Delta\times\mathbb{L}_{1,1}^{3}\Delta$ (0,0,0,0,0,1,0,0,0,2,0,0,4,8,0,0) $\mathbb{L}_{1,0}^{2}\Delta\times\mathbb{L}_{1,1}^{2}\Delta$ (0,0,0,0,0,1,0,0,0,2,0,4,0,8,0,0) $\mathbb{L}_{1,0}\Delta_{2}\times\mathbb{L}_{1,1}^{2}\Delta$ (0,0,0,0,0,1,0,0,0,2,4,8,0,16,0,0) $\mathbb{L}_{6,1}$ (0,0,0,0,0,1,0,0,1,0,1,0,0,0,0,0) $\mathbb{L}_{1,0}\Delta\times\mathbb{L}_{4,1}$ (0,0,0,0,0,1,0,0,1,0,2,0,0,0,0,0) $\mathbb{L}_{1,1}\Delta\times\mathbb{L}_{1,0}^{3}\Delta$ (0,0,0,0,0,1,0,0,2,0,0,4,0,8,0,0) $\mathbb{L}_{1,1}\Delta\times\mathbb{L}_{4,0}$ (0,0,0,0,0,1,0,0,2,0,2,0,0,0,0,0) $\mathbb{L}_{1,0}\Delta^{2}\times\mathbb{L}_{1,1}\Delta$ (0,0,0,0,0,1,0,0,2,0,4,0,0,0,0,0) $\mathbb{L}_{1,0}\Delta\times(\mathbb{L}_{1,1}\times\mathbb{L}_{1,0}\Delta)\Delta$ (0,0,0,0,0,1,0,0,2,0,4,0,8,0,16,0) $\mathbb{L}_{1,1}\Delta\times(\mathbb{L}_{1,0}\times\mathbb{L}_{1,0}\Delta)\Delta$ (0,0,0,0,0,1,0,0,2,0,4,8,0,16,0,0) $\mathbb{L}_{1,1}\Delta\times\mathbb{L}_{1,0}^{2}\Delta_{2}$ (0,0,0,0,0,1,0,2,4,0,8,16,0,32,0,0) $\mathbb{L}_{1,1}\Delta\times\mathbb{L}_{1,0}\Delta_{3}$ (0,0,0,0,0,1,2,4,8,0,16,32,0,64,0,0) $\mathbb{L}_{1,1}\Delta\times\mathbb{L}_{1,1}^{3}\Delta$ (0,0,0,0,1,2,0,0,0,2,0,0,4,0,8,0) $\mathbb{L}_{1,1}^{2}\Delta^{2}$ (0,0,0,0,1,2,0,0,0,2,0,0,4,8,0,0) $\mathbb{L}_{1,1}\Delta\times\mathbb{L}_{4,1}$ (0,0,0,0,1,2,0,0,0,2,1,0,0,0,0,0) $\mathbb{L}_{1,0}\Delta\times\mathbb{L}_{1,1}\Delta^{2}$ (0,0,0,0,1,2,0,0,0,2,4,0,0,0,0,0) $\mathbb{L}_{1,1}\Delta\times(\mathbb{L}_{1,1}\times\mathbb{L}_{1,0}\Delta)\Delta$ (0,0,0,0,1,2,0,0,0,2,4,0,8,0,16,0) $\mathbb{L}_{1,0}\Delta\times(\mathbb{L}_{1,1}\times\mathbb{L}_{1,1}\Delta)\Delta$ (0,0,0,0,1,2,0,0,0,4,8,0,0,0,0,16) $\mathbb{L}_{1,0}\Delta\times\mathbb{L}_{1,1}^{2}\Delta_{2}$ (0,0,0,0,1,2,0,0,4,8,16,0,0,0,0,32) $\mathbb{L}_{1,0}^{2}\Delta\times\mathbb{L}_{1,1}\Delta_{2}$ (0,0,0,0,1,2,0,4,0,0,8,0,0,0,0,16) $\mathbb{L}_{1,0}\Delta_{2}\times\mathbb{L}_{1,1}\Delta_{2}$ (0,0,0,0,1,2,4,8,0,0,16,0,0,0,0,32) $\mathbb{L}_{1,1}^{2}\Delta\times\mathbb{L}_{1,1}\Delta_{2}$ (0,0,0,1,2,4,0,0,0,4,0,0,8,16,0,0) $\mathbb{L}_{1,1}\Delta^{3}$ (0,0,0,1,2,4,0,0,0,4,0,2,0,0,0,0) $\mathbb{L}_{1,1}\Delta\times(\mathbb{L}_{1,1}\times\mathbb{L}_{1,1}\Delta)\Delta$ (0,0,0,1,2,4,0,0,0,4,0,8,0,16,0,0) $\mathbb{L}_{1,1}\Delta\times\mathbb{L}_{1,1}^{2}\Delta_{2}$ (0,0,0,1,2,4,0,0,0,4,8,16,0,32,0,0) $\mathbb{L}_{1,0}\Delta\times\mathbb{L}_{1,1}\Delta_{3}$ (0,0,0,1,2,4,8,0,0,0,0,0,16,32,64) $\mathbb{L}_{1,1}\Delta\times\mathbb{L}_{1,1}\Delta_{3}$ (0,0,1,2,4,8,0,0,0,8,16,32,0,64,0,0) $\mathbb{L}_{1,1}\Delta_{2}^{2}$ (0,0,1,2,4,8,0,0,4,8,16,0,0,0,0,32)
normal quaternionic matrices of order 6, level $>2$
(nondegenerate, nonpythagorean, nonrigid) $\mathbb{L}_{3}\times\mathbb{L}_{1,0}^{2}\Delta$ (1,0,0,0,0,0,0,0,0,1,0,2,0,4,0,0) $\mathbb{L}_{3}\times\mathbb{L}_{1,0}\Delta_{2}$ (1,0,0,0,0,0,0,0,0,1,2,4,0,8,0,0) $\mathbb{L}_{1}\times\mathbb{L}_{1,0}\Delta\times\mathbb{L}_{1,0}^{2}\Delta$ (1,0,0,0,0,0,0,0,0,2,0,0,4,8,0,0) $\mathbb{L}_{1}\times\mathbb{L}_{1,0}\Delta\times\mathbb{L}_{1,0}\Delta_{2}$ (1,0,0,0,0,0,0,0,0,2,4,8,0,16,0,0) $\mathbb{L}_{3}\times\mathbb{L}_{1,1}^{2}\Delta$ (1,0,0,0,0,2,0,0,0,2,0,0,4,1,0,0) $\mathbb{L}_{1}\times\mathbb{L}_{1,0}\Delta\times\mathbb{L}_{1,1}^{2}\Delta$ (1,0,0,0,0,2,0,0,0,2,0,0,4,8,0,0) $\mathbb{L}_{1}\times\mathbb{L}_{1,1}\Delta\times\mathbb{L}_{1,0}^{2}\Delta$ (1,0,0,0,0,2,0,0,0,2,0,4,0,8,0,0) $\mathbb{L}_{1}\times\mathbb{L}_{1,1}\Delta\times\mathbb{L}_{1,0}\Delta_{2}$ (1,0,0,0,0,2,0,0,0,2,4,8,0,16,0,0) $\mathbb{L}_{1,0}\Delta\times(\mathbb{L}_{1,1}^{2}\times\mathbb{L}_{1})\Delta$ (1,0,0,0,0,2,0,0,0,4,0,0,8,16,0,0) $\mathbb{L}_{1,0}^{2}\Delta\times(\mathbb{L}_{1,1}\times\mathbb{L}_{1})\Delta$ (1,0,0,0,0,2,0,0,0,4,0,8,0,16,0,0) $(\mathbb{L}_{1,1}\times\mathbb{L}_{1})\Delta\times\mathbb{L}_{1,0}\Delta^{2}$ (1,0,0,0,0,2,0,0,0,4,8,16,0,32,0,0) $\mathbb{L}_{1,0}\Delta\times\mathbb{L}_{3}\Delta$ (1,0,0,0,0,2,0,0,1,4,8,0,0,0,0,16) $\mathbb{L}_{1}\times\mathbb{L}_{5}$ (1,0,0,0,0,2,0,0,2,0,2,0,0,0,0,0) $\mathbb{L}_{3}^{2}$ (1,0,0,0,0,2,0,0,2,0,3,0,0,0,0,0) $\mathbb{L}_{1}\times\mathbb{L}_{3}\times\mathbb{L}_{1,0}\Delta$ (1,0,0,0,0,2,0,0,2,0,4,0,0,0,0,0) $\mathbb{L}_{1}^{2}\times\mathbb{L}_{4,0}$ (1,0,0,0,0,2,0,0,4,0,4,0,0,0,0,0) $\mathbb{L}_{1}^{2}\times\mathbb{L}_{1,0}\Delta^{2}$ (1,0,0,0,0,2,0,0,4,0,8,0,0,0,0,0) $\mathbb{L}_{1,0}\Delta\times(\mathbb{L}_{1}\times\mathbb{L}_{1,0}\Delta)\Delta$ (1,0,0,0,0,2,0,0,4,0,8,0,16,0,32,0) $\mathbb{L}_{1}\times\mathbb{L}_{1,1}\Delta\times\mathbb{L}_{1,1}^{2}\Delta$ (1,0,0,0,2,4,0,0,0,4,0,0,8,2,0,0) $\mathbb{L}_{1,1}^{2}\Delta\times(\mathbb{L}_{1,1}\times\mathbb{L}_{1})\Delta$ (1,0,0,0,2,4,0,0,0,4,0,0,8,16,0,0) $\mathbb{L}_{1,1}\Delta\times(\mathbb{L}_{1,1}^{2}\times\mathbb{L}_{1})\Delta$ (1,0,0,0,2,4,0,0,0,4,0,8,0,16,0,0) $\mathbb{L}_{1,1}\Delta\times\mathbb{L}_{3}\Delta$ (1,0,0,0,2,4,0,0,0,4,1,8,0,16,0,0) $\mathbb{L}_{1}\times\mathbb{L}_{1,1}\Delta\times\mathbb{L}_{3}$ (1,0,0,0,2,4,0,0,0,4,2,0,0,0,0,0) $\mathbb{L}_{1}^{2}\times\mathbb{L}_{4,1}$ (1,0,0,0,2,4,0,0,0,4,4,0,0,0,0,0) $\mathbb{L}_{1}^{2}\times\mathbb{L}_{1,0}\Delta\times\mathbb{L}_{1,1}\Delta$ (1,0,0,0,2,4,0,0,0,4,8,0,0,0,0,0) $\mathbb{L}_{1,1}\Delta\times(\mathbb{L}_{1}\times\mathbb{L}_{1,0}\Delta)\Delta$ (1,0,0,0,2,4,0,0,0,4,8,16,0,32,0,0) $\mathbb{L}_{3}\times(\mathbb{L}_{1,1}\times\mathbb{L}_{1})\Delta$ (1,0,0,0,2,4,0,0,0,8,2,0,0,0,0,0) $\mathbb{L}_{1}\times\mathbb{L}_{1,0}\Delta\times(\mathbb{L}_{1,1}\times\mathbb{L}_{1})\Delta$ (1,0,0,0,2,4,0,0,0,8,16,0,0,0,0,0) $\mathbb{L}_{1,0}\Delta\times(\mathbb{L}_{1,1}\times\mathbb{L}_{1}^{2})\Delta$ (1,0,0,0,2,4,0,0,0,8,16,0,0,0,0,32) $\mathbb{L}_{3}\times\mathbb{L}_{1,1}\Delta_{2}$ (1,0,0,0,2,4,0,0,2,4,1,0,0,0,0,8) $\mathbb{L}_{1}\times\mathbb{L}_{1,0}\Delta\times\mathbb{L}_{1,1}\Delta_{2}$ (1,0,0,0,2,4,0,0,2,4,8,0,0,0,0,16) $\mathbb{L}_{1,0}\Delta\times(\mathbb{L}_{1}\times\mathbb{L}_{1,1}\Delta)\Delta$ (1,0,0,0,2,4,0,0,2,8,16,0,0,0,0,32) $\mathbb{L}_{1,0}\Delta\times(\mathbb{L}_{1,1}\times\mathbb{L}_{1})\Delta_{2}$ (1,0,0,0,2,4,0,0,8,16,32,0,0,0,0,64) $\mathbb{L}_{1}^{2}\times\mathbb{L}_{1,1}\Delta^{2}$ (1,0,0,2,4,8,0,0,0,8,0,4,0,0,0,0) $\mathbb{L}_{1}\times\mathbb{L}_{1,1}\Delta\times(\mathbb{L}_{1,1}\times\mathbb{L}_{1})\Delta$ (1,0,0,2,4,8,0,0,0,8,0,16,0,0,0,0) $\mathbb{L}_{1,1}\Delta\times(\mathbb{L}_{1,1}\times\mathbb{L}_{1}^{2})\Delta$ (1,0,0,2,4,8,0,0,0,8,0,16,0,32,0,0) $\mathbb{L}_{1}\times\mathbb{L}_{1,1}\Delta\times\mathbb{L}_{1,1}\Delta_{2}$ (1,0,0,2,4,8,0,0,0,8,2,4,0,16,0,0) $\mathbb{L}_{1,1}\Delta\times(\mathbb{L}_{1}\times\mathbb{L}_{1,1}\Delta)\Delta$ (1,0,0,2,4,8,0,0,0,8,2,16,0,32,0,0) $\mathbb{L}_{1,1}\Delta\times(\mathbb{L}_{1,1}\times\mathbb{L}_{1})\Delta_{2}$ (1,0,0,2,4,8,0,0,0,8,16,32,0,64,0,0) $(\mathbb{L}_{1,1}\times\mathbb{L}_{1})\Delta^{2}$ (1,0,0,2,4,8,0,0,0,16,0,32,0,0,8,0) $(\mathbb{L}_{1,1}\times\mathbb{L}_{1})\Delta\times\mathbb{L}_{1,1}\Delta_{2}$ (1,0,0,2,4,8,0,0,0,16,2,4,0,32,0,0) $\mathbb{L}_{1}^{3}\times\mathbb{L}_{3}$ (1,0,0,2,4,8,2,0,0,0,0,0,0,0,0,0) $\mathbb{L}_{3}\times\mathbb{L}_{1}^{2}\Delta$ (1,0,0,2,4,8,2,0,0,0,0,0,0,0,0,16)
[…]